### v4

# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

Abstract.
We establish a quantitative relationship between mixed cohomology classes and the geometric complexity of cohomologically calibrated metric connections with totally skew torsion on product manifolds. We show that the dimension of the off-diagonal curvature subspace of a connection $\nabla^{C}$ is bounded below by the sum of tensor ranks of the mixed Künneth components of its calibration class, minus the dimension of the space of harmonic parallel forms on the first factor. The bound depends only on the mixed class $\omega]_{\mathrm{mixed}}\in H^{3}(M;\mathbb{R})$ and on the space of $g_{1}$-parallel harmonic $2$-forms on $M_{1}$; it is therefore independent of the factor metric $g_{2}$ and of any deformation of $g_{1}$ that preserves parallel $2$-forms, hence it is metric-invariant modulo parallel-form preserving deformations. This provides a computational criterion for geometric complexity and quantifies the interaction between topology and curvature, yielding a quantified version of “forced irreducibility” via the dimension of $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$.
Keywords: Connections with Torsion, Product Manifolds, Holonomy Algebra, De Rham Cohomology, Hodge Theory, Künneth Decomposition, Curvature Tensor
MSC Classification 2020: 53C05, 53C07, 53C29, 58A14, 57R19, 15A69
1. Introduction
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$. We introduce a topological quantity, the mixed tensor rank of the cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that it furnishes, for any metric connection $\nabla^{\mathcal{C}}$, a quantified lower bound for the dimension of the subspace of the holonomy algebra generated by off-diagonal curvature operators with respect to the splitting $V_{1}\oplus V_{2}$. The key point is that, although the definition of “off-diagonal” uses $g$, the lower bound is independent of the factor metric $g_{2}$ and of any deformation of $g_{1}$ that preserves the space of parallel 2-forms, and is explicitly corrected by the dimension of $g_{1}$-parallel harmonic 2-forms on the first factor. Each independent mixed Künneth component in $\omega]$ guarantees a linearly independent off-diagonal curvature direction, modulo the parallel subspace.
2. Cohomologically Calibrated Metric Connections with Skew Torsion
Throughout, $\nabla^{LC}$ denotes the Levi-Civita connection of $g$. We consider metric connections $\nabla^{C}$ with totally skew torsion $T$, i.e.
$\nabla^{C}g=0,$ (2.1)
$T(X,Y,Z):=g(T(X,Y),Z)\in\Omega^{3}(M),$ (2.2)
and write
$\nabla^{C}=\nabla^{LC}+K,$ (2.3)
$K_{XYZ}=\tfrac{1}{2}\,T_{XYZ}.$ (2.4)
Thus $T^{\flat}:=T$ is a 3-form. The curvature $R^{C}$ of $\nabla^{C}$ can be expressed in terms of $R^{LC}$, $\nabla^{LC}T$, and quadratic torsion terms; in abstract index notation,
$R^{C}_{XY}Z=R^{LC}_{XY}Z+\tfrac{1}{2}\big((\nabla^{LC}_{X}T)(Y,Z,\cdot)^{\sharp}-(\nabla^{LC}_{Y}T)(X,Z,\cdot)^{\sharp}\big)+\tfrac{1}{4}\,Q_{T}(X,Y)Z,$ (2.5)
where $Q_{T}$ is bilinear and quadratic in $T$ (see, e.g., [3] or standard torsion-curvature formulas). We shall use (2.5) only qualitatively: mixed components of $T$ produce mixed components of $R^{C}$ through both the $\nabla^{LC}T$ and $T\ast T$ terms.
Definition 2.1 (Metric cohomological calibration).
A metric connection $\nabla^{C}$ with non-parallel ($\nabla^{LC}T\neq 0$) and totally skew torsion $T$, where the cohomology class $\omega]$ is non trivial and defined via the harmonic projection of the torsion three-form, ($T^{\flat}=\omega_{h}+d\alpha+\delta\beta$, so $\omega]:=[\omega_{h}]$), is called cohomologically calibrated. The class is mixed if it lies outside the natural image of $H^{3}(M_{1})\oplus H^{3}(M_{2})$ under the Künneth isomorphism:
$H^{3}(M)\cong\bigoplus_{p+q=3}H^{p}(M_{1})\otimes H^{q}(M_{2}).$ (2.6)
Since $M$ is compact and oriented, Hodge theory applies to $g$: each cohomology class has a unique harmonic representative. We denote by $\omega_{h}$ the harmonic representative of $\omega]$ with respect to $g$.
For a thorough treatment of Hodge theory and cohomology on complex manifolds, see [5].
3. Off-Diagonal Curvature and Holonomy Complexity
Let $P_{i}:T_{p}M\to V_{i}$ be the orthogonal projections of the product splitting.
Definition 3.1 (Off-diagonal curvature).
For a curvature tensor $R$, define its off-diagonal component by
$R_{\mathrm{off}}(X,Y)Z:=P_{1}R(X,Y)(P_{2}Z)+P_{2}R(X,Y)(P_{1}Z).$ (3.1)
Definition 3.2 (Off-diagonal holonomy subspace).
The off-diagonal holonomy subspace at $p\in M$ is
$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\operatorname{Span}_{\mathbb{R}}\{\,R_{\mathrm{off}}(X,Y)Z\;:\;X,Y,Z\in T_{p}M\,\}\ \subseteq\ \mathfrak{hol}_{p}(\nabla^{C}).$ (3.2)
Clearly $\dim\mathfrak{hol}_{p}(\nabla^{C})\geq\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$. Because $\nabla^{C}$ is metric, its holonomy algebra sits in $\mathfrak{so}(T_{p}M)$. If the action were reducible with respect to $V_{1}\oplus V_{2}$, all holonomy endomorphisms would be block-diagonal, forcing $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\{0\}$. Hence nontrivial off-diagonal holonomy implies irreducibility.
4. Mixed Rank of a Degree-$3$ Class
Write the Künneth decomposition of $H^{3}(M)$ as:
$\big(H^{3}(M_{1})\otimes H^{0}(M_{2})\big)\ \oplus\ \big(H^{2}(M_{1})\otimes H^{1}(M_{2})\big)\ \oplus$ (4.1)
$\big(H^{1}(M_{1})\otimes H^{2}(M_{2})\big)\ \oplus\ \big(H^{0}(M_{1})\otimes H^{3}(M_{2})\big).$
Mixed components occur precisely in bidegrees $(2,1)$ and $(1,2)$.
Definition 4.1 (Mixed tensor rank).
Let $\omega]\in H^{3}(M;\mathbb{R})$. Denote by $\pi_{2,1}([\omega])\in H^{2}(M_{1})\otimes H^{1}(M_{2})$ and $\pi_{1,2}([\omega])\in H^{1}(M_{1})\otimes H^{2}(M_{2})$ its mixed Künneth projections. Define
$r_{2,1}=\min\Big\{r\,:\,\pi_{2,1}([\omega])=\sum_{i=1}^{r}\alpha_{i}\otimes\beta_{i},\ \alpha_{i}\in H^{2}(M_{1}),\ \beta_{i}\in H^{1}(M_{2})\Big\},$ (4.2)
and analogously $r_{1,2}$ for $\pi_{1,2}([\omega])$. The mixed rank is
$\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big):=r_{2,1}+r_{1,2}.$ (4.3)
For degree $3$, the tensor rank in each bidegree is well-defined and finite, and can be computed from any decomposition of the harmonic representative $\omega_{h}$ into wedge products of harmonic forms on the factors for the product metric $g$. Although $\omega_{h}$ depends on $g$, the ranks $r_{2,1},r_{1,2}$ are purely topological (indeed, the tensor rank of an element in $H^{2}(M_{1})\otimes H^{1}(M_{2})$ is defined as the minimal number of simple tensors in any decomposition, hence is invariant under change of metric within the product class).
5. Main Result and Proof
Lemma 5.1 (Generic rank of covariant derivatives).
Let $(M_{1},g_{1})$ be a closed oriented smooth Riemannian manifold and $\mathcal{H}^{2}(M_{1})$ the space of harmonic $2$-forms. Fix an $L^{2}$-orthonormal basis $\{\alpha_{i}\}_{i=1}^{r_{2,1}}$ of a subspace of $\mathcal{H}^{2}(M_{1})$ such that the first $s:=\dim\mathcal{P}_{2}$ elements span the parallel subspace $\mathcal{P}_{2}$. Then the set
$\mathcal{U}^{\prime}:=\Bigl\{p_{1}\in M_{1}\;\Big|\;\text{rank}\!\bigl[(\nabla^{g_{1}}\alpha_{i})(p_{1})\bigr]_{i=s+1,\dots,r_{2,1}}=r_{2,1}-s\Bigr\}$
is open and dense in $M_{1}$.
Proof.
Fix smooth local frames of $\Lambda^{2}T^{*}M_{1}$ and $T^{*}M_{1}$ over an open coordinate ball $B\subset M_{1}$. In these frames each $(\nabla^{g_{1}}\alpha_{i})(p_{1})$ becomes a smooth $\mathbb{R}$-valued matrix $A_{i}(p_{1})$ of size $\dim M_{1}\times\binom{\dim M_{1}}{2}$. Form the $(r_{2,1}-s)\times\bigl[\dim M_{1}\cdot\binom{\dim M_{1}}{2}\bigr$] matrix
$\mathbf{A}(p_{1})=\bigl[A_{s+1}(p_{1})^{\top}\!\!,\dots,A_{r_{2,1}}(p_{1})^{\top}\bigr]^{\top}.$
The condition $\text{rank}\mathbf{A}(p_{1})=r_{2,1}-s$ is equivalent to the non-vanishing of at least one $(r_{2,1}-s)\times(r_{2,1}-s)$ minor of $\mathbf{A}(p_{1})$. Each minor is a smooth function on $B$; hence its non-zero locus is open in $B$. Covering $M_{1}$ by such balls we conclude that $\mathcal{U}^{\prime}$ is open. Suppose, by contradiction, that $\mathcal{U}^{\prime}$ is not dense. Then there exists a non-empty open set $V\subset M_{1}$ on which
$\text{rank}\mathbf{A}(p_{1})<r_{2,1}-s\qquad\forall\,p_{1}\in V.$
Thus, for every $p_{1}\in V$, there exists a non-zero vector $\lambda=(\lambda^{s+1},\dots,\lambda^{r_{2,1}})\in\mathbb{R}^{r_{2,1}-s}$ (depending on $p_{1}$) such that
$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(\nabla^{g_{1}}\alpha_{i})(p_{1})=0.$
The smooth section
$\eta:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}\alpha_{i}\in\mathcal{H}^{2}(M_{1})$
then satisfies $\nabla^{g_{1}}\eta\equiv 0$ on $V$. By unique continuation for harmonic forms ([1]), a harmonic form parallel on an open set of any connected component of $M_{1}$ is globally parallel on that component (By Greene–Jacobowitz [4], we approximate $g$ in $C^{\infty}$ by real-analytic metrics $g_{k}$; for each such metric the Hodge-Laplace operator has analytic coefficients, so Aronszajn’s unique-continuation theorem applies to harmonic $2$-forms). Hence, $\eta$ is parallel on the connected component containing $V$, which implies $\eta\in\mathcal{P}_{2}$. This contradicts the linear independence of $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ modulo $\mathcal{P}_{2}$. Therefore $\mathcal{U}^{\prime}$ must be dense in $M_{1}$.
Theorem 5.2 (Topological lower bound for off-diagonal curvature).
Let $M=M_{1}\times M_{2}$ be compact and oriented, $g=g_{1}\oplus g_{2}$ any product metric, and $\nabla^{C}$ any cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]:=[\omega_{h}]\in H^{3}(M;\mathbb{R})$ is a mixed class.
Then for every $p\in M$,
$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\geq\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\!\bigl(\mathcal{P}_{2}\otimes H^{1}(M_{2})\bigr),$
where $\mathcal{P}_{2}\subset H^{2}(M_{1};\mathbb{R})$ denotes the space of harmonic $2$-forms on $M_{1}$ that are parallel with respect to $g_{1}$.
The bound depends only on $\omega]_{\mathrm{mixed}}$ and on the space of $g_{1}$-parallel harmonic $2$-forms; it is therefore invariant under any deformation of $g_{2}$ and under $g_{1}$-deformations that preserve parallel $2$-forms, hence metric-invariant modulo parallel-form preserving deformations.
Proof.
Let $\omega_{h}$ be the $g$-harmonic representative of $\omega]$. Decompose $\omega_{h}$ according to Künneth and Hodge on the factors. For the $(2,1)$-part, write
$\omega_{h}^{2,1}=\sum_{i=1}^{r_{2,1}}\alpha_{i}\wedge\beta_{i},$ (5.1)
with $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$ harmonic and chosen so that the number of terms is minimal. Similarly, write
$\omega_{h}^{1,2}=\sum_{j=1}^{r_{1,2}}\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j},$ (5.2)
with $\tilde{\alpha}_{j}\in\mathcal{H}^{1}(M_{1})$ and $\tilde{\beta}_{j}\in\mathcal{H}^{2}(M_{2})$. The total number of simple mixed tensors equals $r:=r_{2,1}+r_{1,2}=\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)$.
Define the parallel subspace as
$\mathcal{P}_{2}:=\bigl\{\alpha\in\mathcal{H}^{2}(M_{1})\mid\nabla^{g_{1}}\alpha=0\bigr\}\subset H^{2}(M_{1};\mathbb{R})$ (5.3)
and set
$r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)-\dim\!\big(\mathcal{P}_{2}\otimes H^{1}(M_{2})\big).$ (5.4)
We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
Consider the curvature formula (2.5). The mixed contribution of $T=\omega_{h}+\text{(coexact/exact)}$ to $R^{C}$ has two sources. First, the covariant derivative term $\nabla^{LC}T$ produces operators that, when $X,Y$ lie in $V_{1}$ and $Z$ in $V_{2}$ (or vice versa), insert a mixed $3$-form and hence map between $V_{1}$ and $V_{2}$. Second, the quadratic term $Q_{T}$ contains contractions of two copies of $T$ which, for mixed arguments, again produce off-diagonal operators. Pointwise at $p$, for each simple mixed component $\alpha_{i}\wedge\beta_{i}$ one can choose $X,Y\in V_{1}$ so that $\alpha_{i}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\beta_{i}(Z)\neq 0$. Substituting into (2.5) yields an endomorphism with a nontrivial $V_{2}\to V_{1}$ component; dually, for each $\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j}$ one chooses $X\in V_{1}$, $Y,Z\in V_{2}$ to obtain a nontrivial $V_{1}\to V_{2}$ component.
To prove linear independence, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$. By minimality of the tensor-rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists an open-dense subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent. For any such $p$ we pick vectors $X,Y\in T_{p_{1}}M_{1}$ and $Z\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X,Y)|_{p}=\delta_{ik}$ and $\beta_{i}(Z)|_{p}=\delta_{il}$, obtaining $r$ linearly independent endomorphisms $R_{kl}^{\mathrm{off}}(p)\in\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$. This produces at least $r$ linearly independent elements in $\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})$, since curvature endomorphisms generate the holonomy algebra by Ambrose–Singer [2].
For each simple summand $\alpha_{i}\otimes\beta_{i}]$ of the minimal tensor-rank decomposition, choose harmonic representatives $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$. By unique continuation for harmonic forms, their pointwise evaluations are linearly independent on an open-dense subset $\mathcal{U}\subset M$. For every $p=(p_{1},p_{2})\in\mathcal{U}$ the off-diagonal component of the linear part of the curvature,
$R^{\mathrm{off}}(X,Y)Z=\tfrac{1}{2}\big(\nabla^{LC}_{p}T\big)^{\mathrm{off}}(X,Y,Z),$
evaluated on vectors $X,Y\in T_{p_{1}}M_{1}$ and $Z\in T_{p_{2}}M_{2}$, defines a linear map
$\Psi_{p}:\mathrm{span}\{\alpha_{i}\otimes\beta_{i}\}\longrightarrow\mathfrak{so}(T_{p}M)$
whose injectivity follows from the pointwise independence of the evaluations $\{\alpha_{i}(p_{1})\}$ and $\{\beta_{i}(p_{2})\}$: any nontrivial relation among the images would imply a corresponding relation among the tensor products $\alpha_{i}(p_{1})\otimes\beta_{i}(p_{2})$, forcing all coefficients to vanish. Hence $\Psi_{p}$ produces $r$ linearly independent endomorphisms $R^{\mathrm{off}}(X_{i},Y_{i})Z_{i}$ in $\mathfrak{so}(T_{p}M)$.
The linear map
$\Psi_{p}:\mathrm{span}_{\mathbb{R}}\{\alpha_{i}\otimes\beta_{i}\}_{i=1}^{r}\to\mathfrak{so}(T_{p}M)$
defined by
$\Psi_{p}(\alpha_{i}\otimes\beta_{i})(Z):=\tfrac{1}{2}(\nabla^{LC}T)^{\mathrm{off}}(X,Y,Z)$
[…]
The map $\Psi_{p}$ is defined by the linear part of the curvature, which is proportional to $(\nabla^{LC}T)_{\text{off}}$. The condition that a linear combination of these contributions vanishes at a generic point $p=(p_{1},p_{2})$ is:
$\sum_{i}c_{i}\left[(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})\right]_{p}=0$
Since the forms $\alpha_{i}$ and $\beta_{i}$ are linearly independent and the covariant derivatives $\nabla^{g_{1}}$ and $\nabla^{g_{2}}$ act only on their respective factors, the vanishing of the mixed components of this sum implies that the terms must vanish separately.
Specifically, considering the component of type $(2,1)$ (forms with two indices on $M_{1}$ and one on $M_{2}$), the vanishing condition implies:
$\sum_{i}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0$
Since the forms $\beta_{i}$ are linearly independent, this requires that the coefficients of $\beta_{i}$ must vanish:
$\sum_{i}c_{i}(\nabla^{g_{1}}\alpha_{i})=0$
This condition holds on an open dense subset $\mathcal{U}\subset M$ (by Lemma 5.1). Since $\alpha_{i}$ are harmonic forms, the unique continuation principle implies that the linear combination $\sum c_{i}\alpha_{i}$ is a $\nabla^{g_{1}}$-parallel 2-form on $M_{1}$. By definition, the space of such forms is $\mathcal{P}_{2}$.
$\sum_{i}c_{i}\alpha_{i}\in\mathcal{P}_{2}$
This proves that the kernel of the map $\Psi_{p}$ is exactly the subspace generated by the simple tensors whose $\alpha$ component is $\nabla^{g_{1}}$-parallel.
Therefore, the kernel of the map $\Psi_{p}$ is precisely the intersection of the subspace spanned by the minimal decomposition with the space of parallel components:
$\ker\Psi_{p}=\left(\text{span}\{\alpha_{i}\otimes\beta_{i}\}_{i=1}^{r_{2,1}}\right)\cap\left(\mathcal{P}_{2}\otimes H^{1}(M_{2})\right)$
Since $\{\alpha_{i}\otimes\beta_{i}\}$ forms a basis for the mixed component $\pi_{2,1}([\omega])$, and the $\alpha_{i}$ are linearly independent, we can choose the basis such that the first $s=\dim(\mathcal{P}_{2})$ elements of $\{\alpha_{i}\}$ are a basis for $\mathcal{P}_{2}$. Given the linear independence of the $\beta_{i}$, the dimension of the kernel is precisely the number of simple tensors in the minimal decomposition whose $\alpha$-component is parallel. This dimension is $\dim(\ker\Psi_{p})=\dim\left(\pi_{2,1}([\omega])\cap(\mathcal{P}_{2}\otimes H^{1}(M_{2}))\right)$.
This establishes the dimension of the off-diagonal curvature generated by the linear term $(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$ on the open-dense subset $\mathcal{U}^{\prime}$ given by Lemma 5.1.
Consider the one-parameter family of connections
$\nabla^{(s)}=\nabla^{LC}+\frac{s}{2}T.$
Then the off-diagonal curvature satisfies $R_{\text{off}}(s)=s(\nabla^{LC}T)_{\text{off}}+s^{2}Q_{\text{off}}$. Differentiating at $s=0$, we see that the image of $(\nabla^{LC}T)_{\text{off}}$ has dimension at least $r^{\sharp}$. The key step is to ensure that the rank of $R_{\text{off}}(s)$ at $s=1$ is not reduced by the quadratic term $Q_{\text{off}}$.
We consider the one-parameter family of connections $\nabla^{(s)}=\nabla^{LC}+s\frac{1}{2}T$. The off-diagonal curvature is given by $R_{\mathrm{off}}(s)=sA_{1}+s^{2}A_{2}$, where $A_{1}=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ and $A_{2}=Q_{\mathrm{off}}$. On the open-dense set $U$ (defined in Lemma 5.1), the matrix $A_{1}$ has rank $\mathrm{rank}(A_{1})=r^{\sharp}$.
To prove that $\mathrm{rank}(R_{\mathrm{off}}(1))\geq r^{\sharp}$ for the smooth ($C^{\infty}$) metric $g$, we employ the method of analytic approximation and stability of rank (as stated in the Remark 5.6).
By the density theorem of Greene–Jacobowitz [4] and the analyticity of harmonic forms (Morrey–Nirenberg [8]), there exists a sequence of real-analytic metrics $g_{k}\to g$ in the $C^{\infty}$-topology such that $\dim\mathcal{P}_{2}(g_{k})=\dim\mathcal{P}_{2}(g)$ for every $k$; this is possible because the stratum $\{g^{\prime}\mid\dim\mathcal{P}_{2}(g^{\prime})=\dim\mathcal{P}_{2}(g)\}$ is open in the $C^{\infty}$-topology and contains real-analytic metrics densely (Nash–Moser, [9]). Consequently the curvature $R_{\mathrm{off},k}(s)$ associated with $g_{k}$ is real-analytic in $(s,p)$ and the integer $r^{\sharp}$ is the same for every $g_{k}$.
For each fixed $k$, the matrix $R_{\mathrm{off},k}(s)$ is analytic. Since $\mathrm{rank}(A_{1,k})=r^{\sharp}$ on $U$, there exists an $r^{\sharp}\times r^{\sharp}$ minor $M_{k}(s)$ of $R_{\mathrm{off},k}(s)$ such that $M_{k}(s)=s^{r^{\sharp}}D_{k}(s)$, where $D_{k}(0)\neq 0$. By the Identity Principle for Analytic Functions, $D_{k}(s)$ is not identically zero, and its set of zeros is discrete. Thus, $\mathrm{rank}\,R_{\mathrm{off},k}(p)\geq r^{\sharp}$ holds for all $p\in U$ and for all $k$.
The rank function is lower-semicontinuous with respect to $C^{0}$ convergence ([10]). Since $R_{\mathrm{off},k}\to R_{\mathrm{off}}$ uniformly on $M$, the inequality is preserved:
$\mathrm{rank}(R_{\mathrm{off}}(p))\geq\liminf_{k\to\infty}\mathrm{rank}(R_{\mathrm{off},k}(p))\geq r^{\sharp}\quad\text{on a dense subset of }M.$
The extension of the inequality to the entire manifold $M$ follows from the fact that the set $\{p\in M\mid\mathrm{rank}\,R_{\mathrm{off}}(p)\geq r^{\sharp}\}$ is open (by lower-semicontinuity) and contains a dense subset, thus coinciding with $M$.
To extend the inequality
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$
from the dense open subset $\mathcal{U}$ to all of $M$, we use the lower-semicontinuity of the rank of a continuous family of linear maps.
The extension of the inequality $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$ to the entire manifold $M$ follows from the lower-semicontinuity of the rank, as established in [10]. Indeed, by Ambrose–Singer the holonomy algebra is generated by the curvature operators; since the rank of a continuous family of Lie-algebra generators is lower-semicontinuous (Warner [10]), the inequality $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$ extends from the dense open set to all of $M$.
Finally, the lower bound depends only on $\omega]$ and on the space of $g_{1}$-parallel harmonic $2$-forms; it is therefore invariant under any deformation of $g_{2}$ and under any deformation of $g_{1}$ that preserves the space of parallel $2$-forms, hence under any product-metric deformation that leaves $\mathcal{P}_{2}$ unchanged. ∎
Remark 5.3 (Necessity of the Non-Parallel Torsion Condition).
The proof of Theorem 5.1 relies fundamentally on the non-vanishing of the linear term $R^{C}_{\mathrm{off}}\simeq\frac{1}{2}(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$, which is established via the $s\to 0$ limit of the one-parameter family of connections $\nabla^{(s)}$. This dependence on the linear term necessitates the explicit condition $\nabla^{\mathrm{LC}}T\neq 0$ for the proof to be valid.
To illustrate the necessity of this condition, we consider a case where $\nabla^{\mathrm{LC}}T=0$. If $\nabla^{\mathrm{LC}}T=0$, the curvature formula simplifies to $R^{C}=R^{\mathrm{LC}}+\frac{1}{4}Q_{T}$, where $Q_{T}$ is the term quadratic in the torsion $T$. Since $R^{\mathrm{LC}}$ is block-diagonal for a product metric, the off-diagonal curvature is solely determined by $Q_{T}^{\mathrm{off}}$. The proof of Theorem 5.1 provides no guarantee that the quadratic term $Q_{T}^{\mathrm{off}}$ alone is sufficient to generate $r^{\sharp}$ independent directions.
Consider the flat torus $M=T^{3}=T^{2}\times S^{1}$ endowed with the flat product metric $g$. Let $T$ be the constant volume form $T=dx\wedge dy\wedge dz$. This form is harmonic, non-trivial, and its mixed rank is $\text{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})=1$. Since $T$ is constant, we have $\nabla^{\mathrm{LC}}T=0$.
For this specific example, the curvature $R^{C}$ vanishes identically, leading to $\dim(\text{hol}_{p}^{\mathrm{off}}(\nabla^{C}))=0$. The lower bound is calculated as $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2}))=1-1=0$.
The resulting inequality $0\geq 0$ is satisfied, but the bound is trivial. This example demonstrates that when $\nabla^{\mathrm{LC}}T=0$, the topological lower bound $r^{\sharp}$ can be reduced to zero, even for a non-trivial cohomology class ($\omega]\neq 0$). The failure of the bound to provide a non-trivial constraint confirms that the contribution of the quadratic term $Q_{T}^{\mathrm{off}}$ alone is insufficient to guarantee the non-triviality of the topological lower bound, necessitating the explicit condition $\nabla^{\mathrm{LC}}T\neq 0$ for the proof of Theorem 5.1 to hold and to yield a meaningful geometric result.  [^1] The bound is recovered if we consider a non-parallel torsion $T=\omega_{h}+\varepsilon d(\sin(x)dy\wedge dz)$, which maintains the same cohomology class $\omega_{h}]$ but ensures $\nabla^{\mathrm{LC}}T\neq 0$.
Remark 5.4 (Geometric Significance of Invariance).
The invariance of the lower bound under $g_{1}$-deformations that preserve $\mathcal{P}_{2}$ is non-trivial and geometrically significant on manifolds $M_{1}$ with special holonomy (e.g., Calabi-Yau or flat tori). In these cases, the space of parallel forms $\mathcal{P}_{2}$ is non-zero, and the existence of non-homothetic deformations preserving $\mathcal{P}_{2}$ is guaranteed by known results in differential geometry (e.g., deformation theory of special holonomy metrics). This highlights the topological nature of the bound, which is stable under a non-trivial class of metric variations.
Remark 5.5 (Forced irreducibility and quantified off-diagonal holonomy).
Since $\nabla^{C}$ is metric, any non-trivial off-diagonal holonomy implies that $\mathfrak{hol}_{p}(\nabla^{C})$ acts irreducibly on $T_{p}M$. Theorem 5.2 quantifies the “forced irreducibility” principle in terms of the mixed tensor rank, modulo the parallel subspace $\mathcal{P}_{2}\otimes H^{1}(M_{2})$. Specifically, the dimension of the off-diagonal holonomy subspace is bounded below by
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2})),$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$. Irreducibility is forced if and only if $r^{\sharp}>0$, i.e. when the mixed tensor rank exceeds the dimension of the parallel subspace.
Remark 5.6 (Analyticity as a proof device).
The use of real-analytic metrics in the proof of Theorem 5.2 is purely instrumental: analytic approximation (Greene–Jacobowitz) and lower-semicontinuity of matrix rank extend the bound to every $C^{\infty}$ product metric without additional assumptions.
6. Examples and Sharpness
The case $M=S^{2}\times\Sigma_{g}$
Let $\Sigma_{g}$ be a closed oriented surface of genus $g\geq 1$. By Künneth,
$H^{3}(S^{2}\times\Sigma_{g};\mathbb{R})\cong H^{2}(S^{2})\otimes H^{1}(\Sigma_{g}),$ (6.1)
since $H^{1}(S^{2})=0$ and $H^{2}(\Sigma_{g})\cong\mathbb{R}$. The mixed component is therefore isomorphic to $\mathbb{R}^{2g}$. Choosing a harmonic basis $\{\alpha_{j}\}_{j=1}^{2g}$ for $H^{1}(\Sigma_{g})$ and the volume form $\mathrm{vol}_{S^{2}}$ on $S^{2}$, any mixed class has a harmonic representative of the form
$\omega_{h}=\sum_{j=1}^{2g}c_{j}\,\mathrm{vol}_{S^{2}}\wedge\alpha_{j}.$ (6.2)
Thus $\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)=2g$. For each $j$, select $X,Y\in V_{1}$ so that $\mathrm{vol}_{S^{2}}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\alpha_{j}(Z)\neq 0$; the corresponding curvature endomorphisms produce $2g$ independent off-diagonal directions. Hence
$\dim\big(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\big)\ \geq\ 2g-\dim(\mathcal{P}_{2}\otimes H^{1}(\Sigma_{g})),$ (6.3)
where $\mathcal{P}_{2}\subset H^{2}(S^{2};\mathbb{R})$ is the subspace of harmonic parallel $2$-forms on $S^{2}$ (here $\mathcal{P}_{2}=\{0\}$, so the bound is $2g$).
The case $M=S^{2}\times T^{2}$
Here $H^{3}(M)\cong H^{2}(S^{2})\otimes H^{1}(T^{2})\cong\mathbb{R}^{2}$. Writing $\omega_{h}=c_{1}\,\mathrm{vol}_{S^{2}}\wedge dx+c_{2}\,\mathrm{vol}_{S^{2}}\wedge dy$ with $dx,dy$ the harmonic $1$-forms on the torus, one obtains two independent mixed components, hence
$\dim\big(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\big)\ \geq\ 2-\dim(\mathcal{P}_{2}\otimes H^{1}(T^{2})),$ (6.4)
where $\mathcal{P}_{2}=\{0\}$ on $S^{2}$, so the bound is $2$.
7. Concluding Remarks and Perspectives
We have shown that for every compact and orientable product manifold, endowed with any product metric $g$ and for every metric connection satisfying $\nabla^{\mathrm{LC}}T\neq 0$, with totally skew torsion calibrated by a mixed cohomology class, the off-diagonal holonomy dimension admits a quantified lower bound
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2})),$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ and $\mathcal{P}_{2}\subset H^{2}(M_{1};\mathbb{R})$ is the space of harmonic parallel $2$-forms on $M_{1}$. The resulting lower bound is metric-invariant modulo parallel-form preserving deformations, thus providing the first computable obstruction to reducible holonomy on product manifolds. This quantifies how the topological “mixing strength” of $\omega]$ forces geometric “entanglement” across the product splitting, modulo the parallel subspace. The examples above demonstrate sharpness in natural settings where the torsion equals the harmonic representative and the parallel subspace has minimal dimension. It is natural to conjecture that equality holds whenever the torsion is purely harmonic, whereas strict inequality may occur when the coexact or exact parts of $T$ generate additional independent off-diagonal curvature directions or when the parallel subspace has maximal dimension. A systematic classification of such cases is an interesting direction for future work. A further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].
Compactness and Hodge theory. Compactness ensures the existence and uniqueness of harmonic representatives [10], and the vanishing of boundary terms in integrations by parts. Extending the results to noncompact manifolds would require an $L^{2}$ Hodge framework or alternative analytic hypotheses.