### v9

# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

Abstract.
We establish a quantitative relationship between mixed cohomology classes and the geometric complexity of metric connections with totally skew torsion on product manifolds. For any such connection whose non‑parallel torsion 33‑form represents a mixed de Rham class via its harmonic projection, we prove that the dimension of the off‑diagonal holonomy subspace satisfies the lower bound
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
where $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the direct sum of the kernels of the linear obstruction maps $\Psi_{p},\widetilde{\Psi}_{p}$. These kernels are defined intrinsically as the intersections
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\quad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
where $\mathcal{V}_{2,1}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ (respectively $\mathcal{V}_{1,2}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2})$) denotes the linear span of the harmonic representatives of the $(2,1)$- (respectively $(1,2)$-) Künneth component of $\omega]$. The subspaces $\mathcal{V}_{2,1}$ and $\mathcal{V}_{1,2}$ are independent of the choice of a minimal harmonic decomposition. Consequently, the integer $\dim\mathcal{K}$ depends only on the mixed class $\omega]_{\mathrm{mixed}}$ and on the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ of the factor metrics; hence the bound is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata. This yields a computable, metric‑invariant obstruction to reducible holonomy and quantifies the topological “mixing strength” that forces irreducibility across the product splitting.
Keywords: Connections with Torsion, Product Manifolds, Holonomy Algebra, De Rham Cohomology, Hodge Theory, Künneth Decomposition, Curvature Tensor
MSC Classification 2020: 53C05, 53C07, 53C29, 58A14, 57R19, 15A69
1. Introduction
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$. We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
More precisely, we prove
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
where $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the direct sum of the kernels of the natural linear obstruction maps. To define these kernels, let
$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),\qquad\mathcal{V}_{1,2}:=\operatorname{span}\{\widetilde{\alpha}_{j}\otimes\widetilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
where $\{\alpha_{i}\otimes\beta_{i}\}$ (resp. $\{\widetilde{\alpha}_{j}\otimes\widetilde{\beta}_{j}\}$) is any minimal harmonic decomposition of the $(2,1)$- (resp. $(1,2)$-) Künneth component of $\omega]$. Because any two minimal harmonic decompositions span the same subspaces $U\subset\mathcal{H}^{2}(M_{1})$ and $V\subset\mathcal{H}^{1}(M_{2})$, the spaces $\mathcal{V}_{2,1}=U\otimes V$ and $\mathcal{V}_{1,2}$ are well‑defined, i.e., independent of the chosen decomposition. Then
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
The integer $\dim\mathcal{K}$ measures precisely the part of the mixed class that is absorbed by the parallel‑form strata and therefore does not contribute to off‑diagonal curvature. The quantity $r^{\sharp}$ is metric‑invariant under deformations that preserve the parallel‑form strata: it depends only on $\omega]_{\mathrm{mixed}}$ and on the subspaces $\mathcal{P}_{k}(M_{i})$. Consequently, each independent mixed Künneth tensor guarantees a linearly independent off‑diagonal holonomy direction, modulo the parallel forms captured by $\mathcal{K}$.
This yields a computable, purely topological criterion for forced irreducibility of the holonomy representation across the product splitting: irreducibility is forced whenever $r^{\sharp}>0$, i.e. whenever $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})>\dim\mathcal{K}$.
2. Cohomologically Calibrated Metric Connections with Skew Torsion
Throughout, $\nabla^{LC}$ denotes the Levi-Civita connection of $g$. We consider metric connections $\nabla^{C}$ with totally skew torsion $T$, i.e.
$\nabla^{C}g=0,$ (2.1)
$T(X,Y,Z):=g(T(X,Y),Z)\in\Omega^{3}(M),$ (2.2)
and write
$\nabla^{C}=\nabla^{LC}+K,$ (2.3)
$K_{XYZ}=\tfrac{1}{2}\,T_{XYZ}.$ (2.4)
Thus $T^{\flat}:=T$ is a 3-form. The curvature $R^{C}$ of $\nabla^{C}$ can be expressed in terms of $R^{LC}$, $\nabla^{LC}T$, and quadratic torsion terms; in abstract index notation,
$R^{C}_{XY}Z=R^{LC}_{XY}Z+\tfrac{1}{2}\big((\nabla^{LC}_{X}T)(Y,Z,\cdot)^{\sharp}-(\nabla^{LC}_{Y}T)(X,Z,\cdot)^{\sharp}\big)+\tfrac{1}{4}\,Q_{T}(X,Y)Z,$ (2.5)
where $Q_{T}$ is bilinear and quadratic in $T$ (see, e.g., [3] or standard torsion-curvature formulas). We shall use (2.5) only qualitatively: mixed components of $T$ produce mixed components of $R^{C}$ through both the $\nabla^{LC}T$ and $T\ast T$ terms.
Definition 2.1 (Metric cohomological calibration).
A metric connection $\nabla^{C}$ with non-parallel ($\nabla^{LC}T\neq 0$) and totally skew torsion $T$, where the torsion 3-form represents a non trivial mixed cohomology class $\omega]$ via its harmonic projection ($T^{\flat}=\omega_{h}+d\alpha+\delta\beta$, i.e. $\omega]:=[\omega_{h}]$), is called cohomologically calibrated. The class is mixed if it lies outside the natural image of $H^{3}(M_{1})\oplus H^{3}(M_{2})$ under the Künneth isomorphism:
$H^{3}(M)\cong\bigoplus_{p+q=3}H^{p}(M_{1})\otimes H^{q}(M_{2}).$ (2.6)
Since $M$ is compact and oriented, Hodge theory applies to $g$: each cohomology class has a unique harmonic representative. We denote by $\omega_{h}$ the harmonic representative of $\omega]$ with respect to $g$.
For a thorough treatment of Hodge theory and cohomology on complex manifolds, see [5].
3. Off-Diagonal Curvature and Holonomy Complexity
Let $P_{i}:T_{p}M\to V_{i}$ be the orthogonal projections of the product splitting.
Definition 3.1 (Off-diagonal curvature).
For a curvature tensor $R$, define its off-diagonal component by
$R_{\mathrm{off}}(X,Y)Z:=P_{1}R(X,Y)(P_{2}Z)+P_{2}R(X,Y)(P_{1}Z).$ (3.1)
Definition 3.2 (Off-diagonal holonomy subspace).
Let $(M,g)$ be a Riemannian manifold and $\nabla^{C}$ a metric connection on $TM$ with totally skew–symmetric torsion. For a point $p\in M$, we define the off–diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ as the Lie subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$ generated by all off–diagonal components of the curvature tensor and its covariant derivatives at $p$, namely
$\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C}):=\mathrm{Lie}\,\Big\langle\big(\nabla^{k}R^{C}\big)_{p}(X_{1},\ldots,X_{k};Y,Z)^{\mathrm{off}}\;\Big|\;k\geq 0,\;X_{i},Y,Z\in T_{p}M\Big\rangle,$
where $(\cdot)^{\mathrm{off}}$ denotes the off–diagonal projection with respect to the product splitting $T_{p}M=V_{1}\oplus V_{2}$.
By the Ambrose–Singer theorem [2], $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ depends only on the global holonomy algebra of $\nabla^{C}$ and is therefore independent of the choice of local generators and of the pointwise rank of the curvature tensor. In particular, at each point $p\in M$ one has, clearly, $\dim\mathfrak{hol}_{p}(\nabla^{C})\geq\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$. Because $\nabla^{C}$ is metric, its holonomy algebra sits in $\mathfrak{so}(T_{p}M)$. If the action were reducible with respect to $V_{1}\oplus V_{2}$, all holonomy endomorphisms would be block-diagonal, forcing $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\{0\}$. Hence nontrivial off-diagonal holonomy implies irreducibility.
4. Mixed Rank of a Degree-$3$ Class
Write the Künneth decomposition of $H^{3}(M)$ as:
$\big(H^{3}(M_{1})\otimes H^{0}(M_{2})\big)\ \oplus\ \big(H^{2}(M_{1})\otimes H^{1}(M_{2})\big)\ \oplus$ (4.1)
$\big(H^{1}(M_{1})\otimes H^{2}(M_{2})\big)\ \oplus\ \big(H^{0}(M_{1})\otimes H^{3}(M_{2})\big).$
Mixed components occur precisely in bidegrees $(2,1)$ and $(1,2)$.
Definition 4.1 (Mixed tensor rank).
Let $\omega]\in H^{3}(M;\mathbb{R})$. Denote by $\pi_{2,1}([\omega])\in H^{2}(M_{1})\otimes H^{1}(M_{2})$ and $\pi_{1,2}([\omega])\in H^{1}(M_{1})\otimes H^{2}(M_{2})$ its mixed Künneth projections. Define
$r_{2,1}=\min\Big\{r\,:\,\pi_{2,1}([\omega])=\sum_{i=1}^{r}\alpha_{i}\otimes\beta_{i},\ \alpha_{i}\in H^{2}(M_{1}),\ \beta_{i}\in H^{1}(M_{2})\Big\},$ (4.2)
and analogously $r_{1,2}$ for $\pi_{1,2}([\omega])$. The mixed rank is
$\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big):=r_{2,1}+r_{1,2}.$ (4.3)
For degree $3$, the tensor rank in each bidegree is well-defined and finite, and can be computed from any decomposition of the harmonic representative $\omega_{h}$ into wedge products of harmonic forms on the factors for the product metric $g$. Although $\omega_{h}$ depends on $g$, the ranks $r_{2,1},r_{1,2}$ are purely topological (indeed, the tensor rank of an element in $H^{2}(M_{1})\otimes H^{1}(M_{2})$ is defined as the minimal number of simple tensors in any decomposition, hence is invariant under change of metric within the product class).
5. Main Result and Proof
Lemma 5.1 (Generic rank of covariant derivatives).
Let $(M_{1},g_{1})$ be a compact and oriented smooth Riemannian manifold and $\mathcal{H}^{2}(M_{1})$ the space of harmonic $2$-forms. Fix an $L^{2}$-orthonormal basis $\{\alpha_{i}\}_{i=1}^{r_{2,1}}$ of a subspace of $\mathcal{H}^{2}(M_{1})$ such that the first $s:=\dim\mathcal{P}_{2}$ elements span the parallel subspace $\mathcal{P}_{2}$. Then the set
$\mathcal{U}^{\prime}:=\Bigl\{p_{1}\in M_{1}\;\Big|\;\operatorname{rank}\!\bigl[(\nabla^{g_{1}}\alpha_{i})(p_{1})\bigr]_{i=s+1,\dots,r_{2,1}}=r_{2,1}-s\Bigr\}$
is open and dense in $M_{1}$.
Proof.
Fix smooth local frames of $\Lambda^{2}T^{*}M_{1}$ and $T^{*}M_{1}$ over an open coordinate ball $B\subset M_{1}$. In these frames each $(\nabla^{g_{1}}\alpha_{i})(p_{1})$ becomes a smooth $\mathbb{R}$-valued matrix $A_{i}(p_{1})$ of size $\dim M_{1}\times\binom{\dim M_{1}}{2}$. Form the $(r_{2,1}-s)\times\bigl[\dim M_{1}\cdot\binom{\dim M_{1}}{2}\bigr$] matrix
$\mathbf{A}(p_{1})=\bigl[A_{s+1}(p_{1})^{\top}\!,\dots,A_{r_{2,1}}(p_{1})^{\top}\bigr]^{\top}.$
The condition $\operatorname{rank}\mathbf{A}(p_{1})=r_{2,1}-s$ is equivalent to the non‑vanishing of at least one $(r_{2,1}-s)\times(r_{2,1}-s)$ minor of $\mathbf{A}(p_{1})$. Each minor is a smooth function on $B$; hence its non‑zero locus is open in $B$. Covering $M_{1}$ by such balls we conclude that $\mathcal{U}^{\prime}$ is open.
Now suppose, for a contradiction, that $\mathcal{U}^{\prime}$ is not dense. Then there exists a non‑empty open set $V\subset M_{1}$ on which
$\operatorname{rank}\mathbf{A}(p_{1})<r_{2,1}-s\qquad\forall\,p_{1}\in V.$
Since the rank is lower‑semicontinuous, after possibly shrinking $V$ we may assume that the rank is constant, say $\rho<r_{2,1}-s$, on $V$. Consequently, there exists a smooth map $\lambda:V\to\mathbb{R}^{r_{2,1}-s}\setminus\{0\}$ such that
$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{1})(\nabla^{g_{1}}\alpha_{i})(p_{1})=0\qquad\forall p_{1}\in V.$ (5.1)
Now we employ analytic approximation. By the theorem of Greene–Jacobowitz [4], we can approximate the metric $g_{1}$ in the $C^{\infty}$ topology by a sequence of real‑analytic metrics $\{g_{1}^{(k)}\}$. For each $k$, denote by $\alpha_{i}^{(k)}$ the $g_{1}^{(k)}$-harmonic representative of the cohomology class $\alpha_{i}]$. The forms $\alpha_{i}^{(k)}$ converge to $\alpha_{i}$ in $C^{\infty}$ as $k\to\infty$.
Because $\lambda$ is smooth and non‑zero on $V$, we may choose a point $p_{0}\in V$ and a subsequence (still denoted by $k$) such that $\lambda(p_{0})\neq 0$ and
$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})(\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)})(p_{0})=0\qquad\text{for all }k.$ (5.2)
Indeed, (5.1) passes to the limit after extracting a suitable subsequence, since all objects depend smoothly on the metric.
For each fixed $k$, the metric $g_{1}^{(k)}$ is real‑analytic. Define
$\eta^{(k)}:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})\,\alpha_{i}^{(k)}\in\mathcal{H}^{2}(M_{1},g_{1}^{(k)}).$
Equation (5.2) states that $\nabla^{g_{1}^{(k)}}\eta^{(k)}(p_{0})=0$. Because $g_{1}^{(k)}$ is analytic, the coefficients of the Hodge–Laplacian are analytic, and Aronszajn’s unique-continuation theorem [1] applies. Since $\eta^{(k)}$ is harmonic and $g_{1}^{(k)}$ is real-analytic, $\eta^{(k)}$ is also real-analytic. The condition $\nabla_{g_{1}^{(k)}}\eta^{(k)}(p_{0})=0$ implies that $\eta^{(k)}$ is parallel in a neighborhood of $p_{0}$. By the unique-continuation principle for parallel forms on real-analytic manifolds (a consequence of Aronszajn’s theorem applied to the analytic coefficients), $\eta^{(k)}$ must be parallel on the entire connected component.
Passing to the limit $k\to\infty$, we obtain that
$\eta:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})\,\alpha_{i}$
is parallel with respect to the original metric $g_{1}$ on the component containing $p_{0}$. Thus $\eta\in\mathcal{P}_{2}(M_{1})$, contradicting the linear independence of $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ modulo $\mathcal{P}_{2}(M_{1})$.
Therefore $\mathcal{U}^{\prime}$ must be dense in $M_{1}$.
The same reasoning applies verbatim to $\mathcal{H}^{k}(M_{i})$ for any $k$ and for both factors $M_{1}$ and $M_{2}$. ∎
Theorem 5.2 (Topological Lower Bound for Off-Diagonal Curvature).
Let $M=M_{1}\times M_{2}$ be a compact and oriented product manifold with a product metric $g=g_{1}\oplus g_{2}$, and let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]\in H^{3}(M;\mathbb{R})$ is a mixed class.
Let
$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),\qquad\mathcal{V}_{1,2}:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
where $\{\alpha_{i}\otimes\beta_{i}\}$ (resp. $\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}$) is any minimal harmonic decomposition of the $(2,1)$- (resp. $(1,2)$-) Künneth component of $\omega]$. Define
$\ker\Psi_{p}:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
and set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$. Then, for every $p\in M$,
$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\geq r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}.$
The bound depends only on the mixed class $\omega]_{\mathrm{mixed}}$ and on the parallel‑form strata $\mathcal{P}_{k}(M_{i})$; hence it is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.
Proof.
Let $\omega_{h}$ be the $g$-harmonic representative of $\omega]$. Decompose $\omega_{h}$ according to Künneth and Hodge on the factors. For the $(2,1)$-part, write
$\omega_{h}^{2,1}=\sum_{i=1}^{r_{2,1}}\alpha_{i}\wedge\beta_{i},$ (5.3)
with $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$ harmonic and chosen so that the number of terms is minimal. Similarly,
$\omega_{h}^{1,2}=\sum_{j=1}^{r_{1,2}}\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j},$ (5.4)
with $\tilde{\alpha}_{j}\in\mathcal{H}^{1}(M_{1})$ and $\tilde{\beta}_{j}\in\mathcal{H}^{2}(M_{2})$. The total number of simple mixed tensors equals $r:=r_{2,1}+r_{1,2}=\operatorname{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)$.
Set $r^{\sharp}:=r-\dim\mathcal{K}$. We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
Consider the curvature formula (2.5). The mixed contribution of $T=\omega_{h}+\text{(coexact/exact)}$ to $R^{C}$ has two sources. First, the covariant derivative term $\nabla^{LC}T$ produces operators that, when $X,Y$ lie in $V_{1}$ and $Z$ in $V_{2}$ (or vice versa), insert a mixed $3$-form and hence map between $V_{1}$ and $V_{2}$. Second, the quadratic term $Q_{T}$ contains contractions of two copies of $T$ which, for mixed arguments, again produce off‑diagonal operators. Pointwise at $p$, for each simple mixed component $\alpha_{i}\wedge\beta_{i}$ one can choose $X,Y\in V_{1}$ so that $\alpha_{i}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\beta_{i}(Z)\neq 0$. Substituting into (2.5) yields an endomorphism with a nontrivial $V_{2}\to V_{1}$ component; dually, for each $\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j}$ one chooses $X\in V_{1}$, $Y,Z\in V_{2}$ to obtain a nontrivial $V_{1}\to V_{2}$ component.
To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$. By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists an open‑dense subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent. For any such $p$ we pick vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X_{i},Y_{i})|_{p}=1$ and $\beta_{i}(Z_{i})|_{p}=1$.
Define the linear map
$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i}).$ (5.5)
The map is well‑defined and linear; its image consists precisely of the off‑diagonal curvature operators generated by the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ arising from the (2,1)(2,1)‑component of $\omega_{h}$.
If $\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$, then the corresponding off‑diagonal curvature contribution vanishes:
$\frac{1}{2}\sum_{i}c_{i}(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i})=0.$
Using the product rule $\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})$ and the fact that $R^{LC}_{\mathrm{off}}=0$, this condition is equivalent to
$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=0.$ (5.6)
Separating the bidegrees in (5.6) yields the two independent conditions
$\sum_{i=1}^{r_{2,1}}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0,\qquad\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})=0.$ (5.7)
Because the forms $\beta_{i}$ (resp. $\alpha_{i}$) are linearly independent on an open‑dense set, we obtain
$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}=0,\qquad\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}=0$ (5.8)
on that set. By the unique‑continuation principle for harmonic forms (Lemma 5.1), these equalities extend globally, hence
$\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$ (5.9)
Conversely, if (5.9) holds, then $\nabla^{LC}\bigl(\sum_{i}c_{i}\alpha_{i}\wedge\beta_{i}\bigr)=0$, which forces $\Psi_{p}\bigl(\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\bigr)=0$. Therefore $\xi=\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}$ belongs to $\ker\Psi_{p}$ if and only if (5.9) is satisfied.
Now we identify this kernel with the intersection $\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr)$. If $\xi\in\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}\otimes\mathcal{P}_{1}\bigr)$, then $\xi$ can be written as a finite sum $\sum_{k}A_{k}\otimes B_{k}$ with $A_{k}\in\mathcal{P}_{2}$, $B_{k}\in\mathcal{P}_{1}$. For each term, $\nabla^{LC}(A_{k}\wedge B_{k})=0$ because $A_{k}$ and $B_{k}$ are parallel, hence $\Psi_{p}(A_{k}\otimes B_{k})=0$. By linearity, $\Psi_{p}(\xi)=0$, so $\xi\in\ker\Psi_{p}$.
For the converse, assume that $\xi=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$. Equation (5.9) gives
$A:=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad B:=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$
Let $U=\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ and $V=\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$. Set $U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1})$ and $V_{\parallel}=V\cap\mathcal{P}_{1}(M_{2})$. Choose any basis $\{\alpha_{1},\dots,\alpha_{s}\}$ of $U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1})$ and complete it to a basis of $U$ by adding vectors $\{\alpha_{s+1},\dots,\alpha_{\dim U}\}$ whose classes are a basis of the quotient $U/U_{\parallel}$ (hence linearly independent modulo $\mathcal{P}_{2}(M_{1})$). By Lemma 5.1, at a generic point $p_{1}\in M_{1}$ the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{\dim U}$ are linearly independent. Since $A$ is parallel, $\nabla^{g_{1}}A\equiv 0$; evaluating at such a $p_{1}$ yields
$\sum_{i>s}c_{i}\nabla^{g_{1}}\alpha_{i}(p_{1})=0,$
and the linear independence forces $c_{i}=0$ for all $i>s$.
An identical construction applied to $M_{2}$ (choosing a basis of $V_{\parallel}$ and completing it to a basis of $V$ such that the remaining vectors form a basis of the quotient $V/V_{\parallel}$) implies $c_{i}=0$ for all $i>t$.
Hence, the only possibly non-zero coefficients are those with $i\leq\min(s,t)$. For such indices, by construction $\alpha_{i}\in\mathcal{P}_{2}(M_{1})$ and $\beta_{i}\in\mathcal{P}_{1}(M_{2})$. Therefore,
$\xi=\sum_{i=1}^{\min(s,t)}c_{i}\alpha_{i}\otimes\beta_{i}\in\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2}).$
Since $\xi\in\mathcal{V}_{2,1}$ by hypothesis, we have shown
$\ker\Psi_{p}\subseteq\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr).$
The opposite inclusion is immediate: any element of $\mathcal{V}_{2,1}$ that belongs to $\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})$ is a finite sum of simple tensors of parallel forms; consequently its covariant derivative vanishes, and hence it lies in $\ker\Psi_{p}$.
Thus we have proved
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr).$
An identical argument applied to the $(1,2)$-component yields
$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
Now consider the induced map on the quotient
$\overline{\Psi}_{p}:\mathcal{V}_{2,1}/\ker\Psi_{p}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\overline{\Psi}_{p}([\xi]):=\Psi_{p}(\xi).$
The map $\overline{\Psi}_{p}$ is well‑defined and linear. It is injective because $\overline{\Psi}_{p}([\xi])=0$ implies $\Psi_{p}(\xi)=0$, hence $\xi\in\ker\Psi_{p}$ and $\xi]=0$. Consequently, $\overline{\Psi}_{p}$ sends a basis of the quotient to a set of linearly independent off‑diagonal curvature operators, and we obtain
$\dim(\operatorname{im}\Psi_{p})=\dim\bigl(\mathcal{V}_{2,1}/\ker\Psi_{p}\bigr)=r_{2,1}-\dim\ker\Psi_{p}.$ (5.10)
An entirely analogous argument for the (1,2)(1,2)‑component gives
$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr)$ (5.11)
and
$\dim(\operatorname{im}\widetilde{\Psi}_{p})=r_{1,2}-\dim\ker\widetilde{\Psi}_{p}.$ (5.12)
Hence, on the open‑dense set $\mathcal{U}$ where the harmonic evaluations are independent, the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ generates at least
$\bigl(r_{2,1}-\dim\ker\Psi_{p}\bigr)+\bigl(r_{1,2}-\dim\ker\widetilde{\Psi}_{p}\bigr)=r-\dim\mathcal{K}=r^{\sharp}$
linearly independent off‑diagonal curvature operators. By the Ambrose–Singer theorem [2], these operators belong to the holonomy algebra, whence
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}\qquad\text{for all }p\in\mathcal{U}.$
Consider the one-parameter family of connections
$\nabla^{(s)}=\nabla^{LC}+\frac{s}{2}T.$
Then the off-diagonal curvature satisfies $R_{\mathrm{off}}(s)=s(\nabla^{LC}T)_{\mathrm{off}}+s^{2}Q_{\mathrm{off}}$. Differentiating at $s=0$, we see that the image of $(\nabla^{LC}T)_{\mathrm{off}}$ has dimension at least $r^{\sharp}$. The key step is to ensure that the rank of $R_{\mathrm{off}}(s)$ at $s=1$ is not reduced by the quadratic term $Q_{\mathrm{off}}$.
The off-diagonal curvature is given by $R_{\mathrm{off}}(s)=sA_{1}+s^{2}A_{2}$, where $A_{1}=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ and $A_{2}=Q_{\mathrm{off}}$. On the open-dense set $U$ (defined in Lemma 5.1), the matrix $A_{1}$ has rank $\mathrm{rank}(A_{1})=r^{\sharp}$. The rank $r^{\sharp}$ is established by the harmonic component $\omega_{h}$ of $T$. The non-harmonic part $T-\omega_{h}=d\alpha+\delta\beta$ contributes to $A_{1}$, but since the bound is a lower bound, this contribution cannot reduce the rank below $r^{\sharp}$ on the dense set $U$ and is therefore absorbed by the overall estimate $\mathrm{rank}(A_{1})\geq r^{\sharp}$. More formally, a systematic cancellation of the off-diagonal contributions by the non-harmonic part $\eta=d\alpha+\delta\beta$ on an open set would require $\nabla\eta=-\nabla\omega_{h}$ locally. By the Unique Continuation Principle, this identity would extend globally, implying $\omega_{h}]=0$, which contradicts our hypothesis of a non-trivial mixed class. This ensures that the topological information carried by $\omega_{h}$ is rigidly encoded in the curvature.
To prove that $\mathrm{rank}(R_{\mathrm{off}}(1))\geq r^{\sharp}$ for the smooth ($C^{\infty}$) metric $g$, we employ the method of analytic approximation and stability of rank (as stated in Remark 5.6).
By the density theorem of Greene–Jacobowitz [4] and the analyticity of harmonic forms (Morrey–Nirenberg [8]), we can select a sequence of real-analytic product metrics $g_{k}=g_{1,k}\oplus g_{2,k}$ converging to $g$ in the $C^{\infty}$ topology and satisfying
$\dim\mathcal{P}_{k}(M_{i},g_{k})\leq\dim\mathcal{P}_{k}(M_{i},g)\qquad(k=1,2).$
This selection is possible because the dimension function $g^{\prime}\mapsto\dim\mathcal{P}_{k}(M_{i},g^{\prime})$ is upper semicontinuous; consequently, the set $\{g^{\prime}\mid\dim\mathcal{P}_{k}(M_{i},g^{\prime})\leq\dim\mathcal{P}_{k}(M_{i},g)\}$ is open in the $C^{\infty}$ topology. Real-analytic metrics are dense in this open set by Greene–Jacobowitz. For such a sequence we have $\dim\mathcal{K}(g_{k})\leq\dim\mathcal{K}(g)$, and therefore
$r^{\sharp}(g_{k})\;:=\;\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}(g_{k})\;\geq\;r^{\sharp}(g).$
For each fixed $k$, the curvature tensor $R_{\mathrm{off},k}(s)$ of the connection $\nabla^{(s)}=\nabla^{LC}(g_{k})+\frac{s}{2}T$ is real-analytic in $(s,p)$. On the dense open set $U$ furnished by Lemma 5.1 we have $\operatorname{rank}(A_{1,k})=r^{\sharp}(g_{k})$, where $A_{1,k}=\frac{1}{2}(\nabla^{LC}(g_{k})T)_{\mathrm{off}}$. Hence there exists an $r^{\sharp}(g_{k})\times r^{\sharp}(g_{k})$ minor $M_{k}(s)$ of $R_{\mathrm{off},k}(s)$ with $M_{k}(s)=s^{r^{\sharp}(g_{k})}D_{k}(s)$ and $D_{k}(0)\neq 0$. By the Identity Principle for Analytic Functions, $D_{k}(s)$ does not vanish identically and its zero set is discrete; consequently, for each $k$ there exists an open–dense subset $U_{k}\subseteq U$ such that
$\operatorname{rank}R_{\mathrm{off},k}(p)\geq r^{\sharp}(g_{k})\qquad\forall p\in U_{k}.$
Let $W:=\bigcap_{k=1}^{\infty}U_{k}$. Since $M$ is compact, it carries a complete Riemannian metric; in a complete metric space, a countable intersection of open dense subsets is dense. Thus $W$ is dense in $U$, hence dense in $M$. For every $p\in W$ and every $k$ we have
$\operatorname{rank}R_{\mathrm{off},k}(p)\geq r^{\sharp}(g_{k}).$
The rank of a continuous family of linear maps is lower semicontinuous with respect to uniform convergence [9, Warner]. Since $R_{\mathrm{off},k}\to R_{\mathrm{off}}$ uniformly on $M$, for every $p\in W$ we obtain
$\operatorname{rank}R_{\mathrm{off}}(p)\;\geq\;\liminf_{k\to\infty}\operatorname{rank}R_{\mathrm{off},k}(p)\;\geq\;\liminf_{k\to\infty}r^{\sharp}(g_{k})\;\geq\;r^{\sharp}(g).$
The set $\{p\in M\mid\operatorname{rank}R_{\mathrm{off}}(p)\geq r^{\sharp}(g)\}$ is open (by lower semicontinuity) and contains the dense subset $W$; therefore it is all of $M$. Thus $\operatorname{rank}R_{\mathrm{off}}(p)\geq r^{\sharp}(g)$ for every $p\in M$.
Finally, by Definition 3.2 the off‑diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ is the Lie subalgebra generated by the off‑diagonal curvature operators and their covariant derivatives at $p$. In particular, it contains the linear span of the off‑diagonal curvature operators $R_{\mathrm{off}}(p)$. Therefore,
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;\operatorname{rank}R_{\mathrm{off}}(p)\;\geq\;r^{\sharp}(g)\qquad\forall p\in M,$
which is the desired bound. The lower bound depends only on $\omega]$ and on the spaces of $g_{1}$‑parallel harmonic 22‑forms and 11‑forms, and of $g_{2}$‑parallel harmonic 11‑forms and 22‑forms; it is therefore invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these parallel‑form strata, hence under any product‑metric deformation that leaves $\mathcal{P}_{k}(M_{i})$ unchanged for $k=1,2$. ∎
Remark 5.3 (Necessity of the Non‑Parallel Torsion Condition).
Theorem 5.2 yields a non‑trivial lower bound only when the linear term
$\tfrac{1}{2}(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$
is non‑zero. If instead $\nabla^{\mathrm{LC}}T=0$, the curvature reduces to
$R^{C}=R^{\mathrm{LC}}+\tfrac{1}{4}Q_{T},$
and there is no guarantee that the purely quadratic off‑diagonal part $(Q_{T})_{\mathrm{off}}$ generates $r^{\sharp}$ independent directions.
Consider for instance the flat torus $M=T^{2}\times S^{1}$ with the constant volume form $T=dx\wedge dy\wedge dz$, which is harmonic and satisfies
$\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})=1.$
Since $T$ is parallel, $\nabla^{\mathrm{LC}}T=0$; hence every simple tensor in any minimal harmonic decomposition is completely parallel. In the notation of Theorem 5.2 we have
$\mathcal{V}_{2,1}=\operatorname{span}\{dx\wedge dy\otimes dz\},\qquad\mathcal{P}_{2}(T^{2})=\operatorname{span}\{dx\wedge dy\},\qquad\mathcal{P}_{1}(S^{1})=\operatorname{span}\{dz\},$
so that
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(T^{2})\otimes\mathcal{P}_{1}(S^{1})\bigr)=\mathcal{V}_{2,1},\qquad\ker\widetilde{\Psi}_{p}=0.$
Thus $\dim\mathcal{K}=1$ and consequently
$r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})-\dim\mathcal{K}=1-1=0.$
A direct computation gives $R^{C}=0$, hence
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})=0,\qquad r^{\sharp}=0.$
The bound $0\geq 0$ is saturated but trivial, illustrating that without the non‑parallel torsion condition the theorem cannot force irreducibility.
Remark 5.4 (Geometric Significance of Invariance).
The invariance of the lower bound $r^{\sharp}$ is a non‑trivial and geometrically significant property. The bound depends only on two pieces of data:
  • the mixed class $\omega]_{\mathrm{mixed}}$,
  • the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ of the factor metrics.
Explicitly,
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
where $\mathcal{V}_{2,1}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ (resp. $\mathcal{V}_{1,2}$) denotes the linear span of the harmonic representatives of the $(2,1)$- (resp. $(1,2)$-) Künneth component of $\omega]$.
Consequently, $r^{\sharp}$ is invariant under any deformation of the product metric $g=g_{1}\oplus g_{2}$ that preserves the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ ($k=1,2$). Such deformations include, for instance, any variation of $g_{i}$ inside the connected component of the moduli space of metrics with a fixed set of parallel forms.
This stability is especially relevant when at least one factor $M_{i}$ possesses special holonomy (e.g., Calabi–Yau manifolds or flat tori), where the spaces $\mathcal{P}^{k}(M_{i})$ are non‑trivial and the metric $g_{i}$ admits large families of deformations that leave $\mathcal{P}^{k}(M_{i})$ unchanged.
Thus, $r^{\sharp}$ quantifies the part of the topological “mixing strength” of $\omega]$ that cannot be “absorbed” by metric deformations respecting the parallel‑form strata; it captures a genuinely topological obstruction to reducible holonomy.
Remark 5.5 (Forced irreducibility and quantified off‑diagonal holonomy).
Since $\nabla^{C}$ is metric, any non‑trivial off‑diagonal holonomy implies that $\mathfrak{hol}_{p}(\nabla^{C})$ acts irreducibly on $T_{p}M$. Theorem 5.2 quantifies this “forced irreducibility” principle in terms of the mixed tensor rank, modulo the obstruction space
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},$
where, in the notation of the theorem,
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
Specifically, the dimension of the off‑diagonal holonomy subspace satisfies
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
and irreducibility is forced precisely when the right‑hand side is positive, i.e. when
$\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})>\dim\mathcal{K}.$
In other words, a non‑zero value of the corrected rank $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}$ guarantees that the holonomy representation cannot split according to the product decomposition $T_{p}M=V_{1}\oplus V_{2}$; the connection is geometrically “entangled” across the two factors.
Remark 5.6 (Analyticity as a tool of proof).
The use of real‑analytic metrics in the proof of Theorem 5.2 is purely instrumental. By the density theorem of Greene–Jacobowitz [4] and the upper semicontinuity of the dimension of the space of parallel forms, one can approximate any smooth product metric $g=g_{1}\oplus g_{2}$ in the $C^{\infty}$-topology by a sequence of real‑analytic product metrics $\{g^{(k)}\}$ satisfying $\dim\mathcal{P}_{\ell}(M_{i},g^{(k)})\leq\dim\mathcal{P}_{\ell}(M_{i},g)$ for $\ell=1,2$.
For each analytic metric $g^{(k)}$ the curvature tensor $R_{\mathrm{off},k}(s)$ depends analytically on the deformation parameter $s$ and on the point $p\in M$; consequently the rank of the off‑diagonal curvature is lower‑semicontinuous and can be controlled by the behaviour of analytic minors (Identity Principle).
Passing to the limit $k\to\infty$ via the lower‑semicontinuity of matrix rank with respect to uniform convergence [9] yields the desired bound for the original $C^{\infty}$ metric $g$. Thus the analytic approximation does not impose any restriction on the result—it merely provides a convenient technical tool for extending the estimate from a generic dense subset to the whole manifold.
6. Illustrative Examples
The case $\Sigma_{g}\times T^{2}$ (Example of a non-trivial lower bound)
Fix $g\geq 2$ and endow $\Sigma_{g}$ with any hyperbolic metric and $T^{2}$ with the flat metric. Harmonic spaces:
$H^{1}(T^{2})=\operatorname{span}\{dx,dy\},\qquad H^{2}(\Sigma_{g})=\operatorname{span}\{\operatorname{vol}_{\Sigma}\},$
and choose a non-zero harmonic $1$-form $\alpha\in H^{1}(\Sigma_{g})$ (such an $\alpha$ exists because $b_{1}(\Sigma_{g})=2g>0$).
For the hyperbolic metric on $\Sigma_{g}$ we have
$\mathcal{P}_{1}(\Sigma_{g})=0,\qquad\mathcal{P}_{2}(\Sigma_{g})=\operatorname{span}\{\operatorname{vol}_{\Sigma}\}\cong\mathbb{R},$
since the volume form is harmonic and parallel.
For the flat metric on $T^{2}$ all harmonic forms are parallel, hence
$\mathcal{P}_{1}(T^{2})=H^{1}(T^{2})\cong\mathbb{R}^{2},\qquad\mathcal{P}_{2}(T^{2})=H^{2}(T^{2})\cong\mathbb{R}.$
Consider the mixed cohomology class
$\omega]=\underbrace{\operatorname{vol}_{\Sigma}\otimes dx}_{\in H^{2}(\Sigma_{g})\otimes H^{1}(T^{2})}\;+\;\underbrace{\alpha\otimes\operatorname{vol}_{T^{2}}}_{\in H^{1}(\Sigma_{g})\otimes H^{2}(T^{2})}\;\in\;H^{3}(\Sigma_{g}\times T^{2}).$
Both Künneth components are already simple tensors, therefore their minimal tensor ranks are
$r_{2,1}=1,\qquad r_{1,2}=1,$
and the mixed rank equals
$\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)=r_{2,1}+r_{1,2}=2.$
The obstruction kernels are computed as follows. For the $(2,1)$-part $\pi_{2,1}([\omega])=\operatorname{vol}_{\Sigma}\otimes dx$,
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr)=\operatorname{span}\{\operatorname{vol}_{\Sigma}\otimes dx\},$
because $\operatorname{vol}_{\Sigma}\in\mathcal{P}_{2}(\Sigma_{g})$ and $dx\in\mathcal{P}_{1}(T^{2})$. Hence $\dim\ker\Psi_{p}=1$.
For the $(1,2)$-part $\pi_{1,2}([\omega])=\alpha\otimes\operatorname{vol}_{T^{2}}$,
$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr)=\{0\},$
because $\mathcal{P}_{1}(M_{1})=\{0\}$ and $\mathcal{P}_{2}(M_{2})=\operatorname{span}\{\operatorname{vol}_{T^{2}}\}$, hence $\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})=\{0\}$. Hence $\dim\ker\widetilde{\Psi}_{p}=0$. Thus
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\dim\mathcal{K}=1+0=1,$
and the corrected rank is
$r^{\sharp}=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}=2-1=1.$
Theorem 5.2 therefore yields the non‑trivial lower bound
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;1.$
In particular, any cohomologically calibrated metric connection with non‑parallel torsion representing $\omega]$ necessarily has irreducible holonomy across the product splitting $\Sigma_{g}\times T^{2}$.
The case $T^{2}\times T^{2}$ (Example of a trivial lower bound)
Let $M=T^{2}\times T^{2}$ endowed with the flat product metric $g=g_{1}\oplus g_{2}$. Fix harmonic bases
$H^{1}(T^{2})=\operatorname{span}\{dx,dy\},\quad H^{2}(T^{2})=\operatorname{span}\{vol\}.$
For the flat metric we have
$\mathcal{P}_{1}(T^{2})=H^{1}(T^{2})\cong\mathbb{R}^{2},\qquad\mathcal{P}_{2}(T^{2})=H^{2}(T^{2})\cong\mathbb{R}.$
The mixed Künneth components are
$H^{2}(T^{2})\otimes H^{1}(T^{2})\cong\mathbb{R}^{2},\quad H^{1}(T^{2})\otimes H^{2}(T^{2})\cong\mathbb{R}^{2}.$
Consider the single cohomology class
$\omega]=\underbrace{vol\otimes dx}_{(2,1)}+\underbrace{dy\otimes vol}_{(1,2)}.$
Its minimal tensor-rank decomposition is
$\pi_{2,1}([\omega])=vol\otimes dx\quad(\text{rank }1),\qquad\pi_{1,2}([\omega])=dy\otimes vol\quad(\text{rank }1).$
Hence
$\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1+1=2.$
Since $vol\in\mathcal{P}_{2}$ and $dx,dy\in\mathcal{P}_{1}$, both simple tensors lie in the obstruction spaces
$\ker\Psi_{p}=\operatorname{span}\{vol\otimes dx\},\qquad\ker\tilde{\Psi}_{p}=\operatorname{span}\{dy\otimes vol\},$
so $\dim\mathcal{K}=2$. Theorem 5.2 gives
$\dim\mathfrak{hol}^{\text{off}}_{p}(\nabla^{C})\geq 2-2=0.$
7. Significance and Applications to Physics and STEM Disciplines
Connections with torsion and product structures appear naturally in a wide range of mathematical and physical models. In string theory, supergravity, and related frameworks, metric connections with totally skew-symmetric torsion play a central role in flux compactifications, generalized geometries, and effective field theories on product manifolds such as $M_{4}\times X$. In these settings, the holonomy of the canonical connection governs the coupling between the external (space-time) and internal (compactified) factors, and irreducibility prevents the decoupling of physical degrees of freedom.
Our lower bound provides an efficient mechanism for establishing such irreducibility directly from global invariants - particularly the mixed tensor rank and the corrected quantity $r^{\sharp}$ - without computing curvature tensors or higher covariant derivatives. This is especially relevant in high-dimensional backgrounds where explicit curvature calculations are notoriously difficult or mathematically intractable.
A similar phenomenon arises in gauge theory, condensed matter, and materials science, where product manifolds model layered, composite, or multi-component systems. The holonomy algebra encodes the geometric interactions between these components; detecting whether such interactions are essential (irreducible) or merely apparent becomes a fundamental structural question. The criterion developed in this paper identifies irreducibility based solely on the existence and non-parallelity of mixed differential forms, providing a geometric diagnostic that is robust under perturbations of the underlying metric and torsion.
Applications also extend to geometric analysis and computational modeling. In multi-scale PDE systems defined on product domain (such as transport-diffusion equations, elasticity models, or geometric flows) the holonomy of a natural geometric connection determines how different subsystems interact and whether the governing equations admit a decoupled formulation. Our lower bound offers a rigorous obstruction to such decoupling, enabling researchers to understand when full coupling is intrinsic to the geometry. This is particularly useful in numerical simulations, where direct computation of curvatures is expensive and unstable, while global invariants such as parallel-form spaces are straightforward to detect algorithmically.
Overall, the ability to certify irreducible holonomy through topological or geometric invariants provides a powerful tool across physics, geometry, and applied sciences. It eliminates the need for explicit curvature computations and transforms a classically analytic problem into a global geometric one, opening the door to applications in any setting where torsion, product structures, and holonomy play a foundational role.
8. Concluding Remarks and Perspectives
We have shown that for every compact and orientable product manifold, endowed with any product metric $g$ and for every metric connection satisfying $\nabla^{\mathrm{LC}}T\neq 0$, with 3-form torsion representing a mixed cohomology class via its harmonic projection $\omega_{h}$, the off-diagonal holonomy dimension admits a quantified lower bound
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K},$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\tilde{\Psi}_{p}$ is the sum of the kernels of the linear maps $\Psi_{p},\tilde{\Psi}_{p}$ that encode the obstruction to off-diagonal curvature. The resulting lower bound is metric-invariant modulo parallel-form preserving deformations, thus providing the first computable obstruction to reducible holonomy on product manifolds. This quantifies how the topological “mixing strength” of $\omega]$ forces geometric “entanglement” across the product splitting, modulo the parallel subspace contained within the mixed class. The examples above illustrate the computation of $r^{\sharp}$ in simple geometric settings. When the torsion is purely harmonic and the parallel-form spaces are minimal, the bound is often saturated in explicit constructions (e.g., on flat tori with parallel torsion). It is natural to conjecture that equality holds generically for purely harmonic torsion, while strict inequality may occur when the coexact or exact parts of $T$ generate additional off-diagonal curvature directions, or when the parallel-form spaces are larger. A systematic classification of the cases where the bound is sharp remains an interesting direction for future work.
A further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].
Compactness and Hodge theory. Compactness ensures the existence and uniqueness of harmonic representatives [9], and the vanishing of boundary terms in integrations by parts. Extending the results to noncompact manifolds would require an $L^{2}$ Hodge framework or alternative analytic hypotheses.