### v10

# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

Abstract.
We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on admissible product manifolds: either both factors are compact oriented surfaces, or $\dim M=4$ with a $(3,1)$ splitting and the mixed class has tensor rank $1$. For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
with $\mathcal{K}$ an intrinsically defined obstruction space. The bound is a topological invariant under metric deformations preserving the parallel‑form strata and provides a sharp obstruction to reducible holonomy. Counterexamples show the hypothesis is optimal.
Keywords: Connections with Torsion, Product Manifolds, Holonomy Algebra, De Rham Cohomology, Hodge Theory, Künneth Decomposition, Curvature Tensor
MSC Classification 2020: 53C05, 53C07, 53C29, 58A14, 57R19, 15A69
1. Introduction
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$. We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
A subtle but crucial point is that the subspaces spanned by a minimal harmonic decomposition are uniquely determined only under topological restrictions on the factors. Consequently, our main result holds for admissible product manifolds (Definition 2.2), i.e.:
  • (A)
    both factors are compact oriented surfaces;  or
  • (B)
    $\dim M=4$ with a $(3,1)$ splitting and the mixed class has total tensor rank $1$.
For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
where the obstruction space $\mathcal{K}$ is intrinsically defined via intersections of the mixed factor spaces $\mathcal{V}_{2,1}$, $\mathcal{V}_{1,2}$ with the parallel‑form strata $\mathcal{P}_{k}(M_{i})$. The bound is invariant under metric deformations preserving these strata.
When $r^{\sharp}>0$ the holonomy representation is forced to be irreducible across the product splitting. The examples in Section 6 illustrate both non‑trivial ($\Sigma_{g}\times T^{2}$) and trivial ($T^{2}\times T^{2}$) cases, and Remark 2.4 shows by explicit counterexample that without the admissibility hypothesis the bound fails to be well‑defined.
2. Cohomologically Calibrated Metric Connections with Skew Torsion
Throughout, $\nabla^{LC}$ denotes the Levi-Civita connection of $g$. We consider metric connections $\nabla^{C}$ with totally skew torsion $T$, i.e.
$\nabla^{C}g=0,$ (2.1)
$T(X,Y,Z):=g(T(X,Y),Z)\in\Omega^{3}(M),$ (2.2)
and write
$\nabla^{C}=\nabla^{LC}+K,$ (2.3)
$K_{XYZ}=\tfrac{1}{2}\,T_{XYZ}.$ (2.4)
Thus $T^{\flat}:=T$ is a 3-form. The curvature $R^{C}$ of $\nabla^{C}$ can be expressed in terms of $R^{LC}$, $\nabla^{LC}T$, and quadratic torsion terms; in abstract index notation,
$R^{C}_{XY}Z=R^{LC}_{XY}Z+\tfrac{1}{2}\big((\nabla^{LC}_{X}T)(Y,Z,\cdot)^{\sharp}-(\nabla^{LC}_{Y}T)(X,Z,\cdot)^{\sharp}\big)+\tfrac{1}{4}\,Q_{T}(X,Y)Z,$ (2.5)
where $Q_{T}$ is bilinear and quadratic in $T$ (see, e.g., [3] or standard torsion-curvature formulas). We shall use (2.5) only qualitatively: mixed components of $T$ produce mixed components of $R^{C}$ through both the $\nabla^{LC}T$ and $T\ast T$ terms.
Definition 2.1 (Metric cohomological calibration).
A metric connection $\nabla^{C}$ with non-parallel ($\nabla^{LC}T\neq 0$) and totally skew torsion $T$, where the torsion 3-form represents a non trivial mixed cohomology class $\omega]$ via its harmonic projection ($T^{\flat}=\omega_{h}+d\alpha+\delta\beta$, i.e. $\omega]:=[\omega_{h}]$), is called cohomologically calibrated. The class is mixed if it lies outside the natural image of $H^{3}(M_{1})\oplus H^{3}(M_{2})$ under the Künneth isomorphism:
$H^{3}(M)\cong\bigoplus_{p+q=3}H^{p}(M_{1})\otimes H^{q}(M_{2}).$ (2.6)
Since $M$ is compact and oriented, Hodge theory applies to $g$: each cohomology class has a unique harmonic representative. We denote by $\omega_{h}$ the harmonic representative of $\omega]$ with respect to $g$.
For a thorough treatment of Hodge theory and cohomology on complex manifolds, see [5].
2.1. Intrinsic Definition of Mixed Cohomology Subspaces
We now address a subtle but crucial point: the subspaces spanned by the factors in a minimal harmonic decomposition are not uniquely determined in general. To obtain a well–defined geometric invariant we must impose conditions on the topology of the factors.
Definition 2.2 (Admissible Product Manifolds).
A compact oriented product manifold $M=M_{1}\times M_{2}$ is called admissible if it satisfies one of the following:
  • (A)
    both factors $M_{1},M_{2}$ are compact oriented surfaces;
  • (B)
    $\dim M=4$ with $(\dim M_{1},\dim M_{2})=(3,1)$ or $(1,3)$, and the mixed class $\omega]$ has tensor rank $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$.
Lemma 2.3 (Uniqueness of Mixed Subspaces).
Let $M$ be an admissible product manifold and $\omega]\in H^{3}(M;\mathbb{R})$ a non–trivial mixed class. Then the subspaces
$\mathcal{V}_{2,1}$ $\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),$ (2.7)
$\mathcal{V}_{1,2}$ $\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2})$ (2.8)
spanned by any minimal harmonic decomposition of the mixed Künneth components of $\omega]$ are intrinsically defined—i.e. they depend only on $\omega]$ and not on the particular minimal decomposition chosen.
Proof.
We treat the two admissible cases separately.
Case (A): $\dim M_{i}=2$. For a compact oriented surface $F^{2}$ we have $b_{2}(F)=1$ and $b_{1}(F)=2g(F)$. Let $\operatorname{vol}_{F}$ be the unique (up to scale) $g_{F}$–harmonic 22–form. Then
$\mathcal{H}^{2}(F)=\mathbb{R}\cdot\operatorname{vol}_{F}.$
Consequently any non–zero $T\in\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ has rank one and can be written uniquely as
$T=\operatorname{vol}_{M_{1}}\otimes\beta,\qquad\beta\in\mathcal{H}^{1}(M_{2}).$
Indeed, if $\operatorname{vol}_{M_{1}}\otimes\beta=\operatorname{vol}_{M_{1}}\otimes\beta^{\prime}$, then $\operatorname{vol}_{M_{1}}\otimes(\beta-\beta^{\prime})=0$ and, because $\operatorname{vol}_{M_{1}}$ is non–zero as an element of $\mathcal{H}^{2}(M_{1})$, we must have $\beta=\beta^{\prime}$. Thus the one–dimensional subspace $\operatorname{span}\{\beta\}$ is uniquely determined by $T$, and
$\mathcal{V}_{2,1}=\mathcal{H}^{2}(M_{1})\otimes\operatorname{span}\{\beta\}$
is intrinsically defined. The same argument applied symmetrically yields the intrinsic definition of $\mathcal{V}_{1,2}$.
Case (B): $(3,1)$ or $(1,3)$ with $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$. Assume $(\dim M_{1},\dim M_{2})=(3,1)$ (the other case is symmetric). Then $b_{1}(M_{2})=1$ and $H^{2}(M_{2})=0$. Let $\gamma$ be a fixed basis of $\mathcal{H}^{1}(M_{2})$. Any $T\in\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ has rank at most one and can be written uniquely as
$T=\alpha\otimes\gamma,\qquad\alpha\in\mathcal{H}^{2}(M_{1}).$
Uniqueness follows as in Case (A) because $\gamma$ is non–zero. Hence $\mathcal{V}_{2,1}=\operatorname{span}\{\alpha\}\otimes\mathcal{H}^{1}(M_{2})$ depends only on $\omega]$. The component $(1,2)$ is trivial since $H^{2}(M_{2})=0$, so $\mathcal{V}_{1,2}=\{0\}$ is obviously well defined. ∎
Remark 2.4 (Failure of invariance in general).
Invariance fails without the admissibility hypothesis. For instance, on $M=T^{3}\times T^{3}$ we have $b_{2}(T^{3})=3$ and $b_{1}(T^{3})=3$. The mixed class
$\omega]=(dx\wedge dy)\otimes dz+(dx\wedge dz)\otimes dy\in H^{2}(T^{3})\otimes H^{1}(T^{3})$
has rank $2$. It admits two distinct minimal decompositions
$\omega_{h}^{2,1}$ $=(dx\wedge dy)\otimes dz+(dx\wedge dz)\otimes dy$
$=\tfrac{1}{2}\big[(dx\wedge dy+dx\wedge dz)\otimes(dz+dy)$
$\qquad\quad+(dx\wedge dy-dx\wedge dz)\otimes(dz-dy)\big],$
which generate different subspaces $\mathcal{V}_{2,1}$ and $\mathcal{V}_{2,1}^{\prime}$. Consequently the integer $\dim\big(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2}))\big)$ depends on the choice, contradicting the metric–invariance.
2.2. Mixed obstruction spaces
Having established the intrinsic nature of the mixed factor spaces in Lemma 2.3 for admissible product manifolds (Definition 2.2), we now define the obstruction kernels.
Definition 2.5 (Mixed obstruction spaces).
For an admissible product manifold $M$ and a mixed class $\omega]$ we define
$\mathcal{V}_{2,1}$ $:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),$
$\mathcal{V}_{1,2}$ $:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
where the spans are taken over any minimal harmonic decomposition of the mixed Künneth components of $\omega]$; by Lemma 2.3 these spans are independent of the particular choice.
The mixed obstruction spaces are the intersections
$\ker\Psi_{p}$ $:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),$
$\ker\widetilde{\Psi}_{p}$ $:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
and we set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$.
3. Off-Diagonal Curvature and Holonomy Complexity
Let $P_{i}:T_{p}M\to V_{i}$ be the orthogonal projections of the product splitting.
Definition 3.1 (Off-diagonal curvature).
For a curvature tensor $R$, define its off-diagonal component by
$R_{\mathrm{off}}(X,Y)Z:=P_{1}R(X,Y)(P_{2}Z)+P_{2}R(X,Y)(P_{1}Z).$ (3.1)
Definition 3.2 (Off-diagonal holonomy algebra).
Let $(M,g)$ be a Riemannian product manifold with orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$, and let $\nabla^{C}$ be a metric connection on $TM$ with totally skew-symmetric torsion. Let $\Pi_{\mathrm{off}}:\mathfrak{so}(T_{p}M)\to\mathfrak{so}(T_{p}M)$ be the linear projection onto off-diagonal endomorphisms:
$\Pi_{\mathrm{off}}(A)(Z):=P_{1}A(P_{2}Z)+P_{2}A(P_{1}Z),$
where $P_{i}:T_{p}M\to V_{i}$ are the orthogonal projections.
Let $\mathcal{G}_{p}:=\{\nabla^{k}R^{C}|_{p}(X_{1},\dots,X_{k};Y,Z)\mid k\geq 0,X_{i},Y,Z\in T_{p}M\}$ be the set of Ambrose–Singer curvature generators of $\mathfrak{hol}_{p}(\nabla^{C})$. Define the off-diagonal curvature span
$\mathcal{S}_{p}^{\mathrm{off}}:=\operatorname{Span}_{\mathbb{R}}\{\Pi_{\mathrm{off}}(A)\mid A\in\mathcal{G}_{p}\}\subset\mathfrak{so}(T_{p}M).$
The off-diagonal holonomy algebra is the Lie subalgebra obtained by intersecting $\mathfrak{hol}_{p}(\nabla^{C})$ with the Lie algebra generated by $\mathcal{S}_{p}^{\mathrm{off}}$ in $\mathfrak{so}(T_{p}M)$:
$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\bigl\langle\mathcal{S}_{p}^{\mathrm{off}}\bigr\rangle_{\mathrm{Lie}}\cap\mathfrak{hol}_{p}(\nabla^{C})\subset\mathfrak{hol}_{p}(\nabla^{C}).$
Here $\langle\cdot\rangle_{\mathrm{Lie}}$ denotes the Lie algebra generated in $\mathfrak{so}(T_{p}M)$, and the intersection ensures that $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ is canonically a subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$. Equivalently, it is the smallest Lie subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$ whose off-diagonal components contain $\mathcal{S}_{p}^{\mathrm{off}}$.
By the Ambrose–Singer theorem [2], $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ depends only on the global holonomy algebra of $\nabla^{C}$ and is therefore independent of the choice of local generators and of the pointwise rank of the curvature tensor. In particular, at each point $p\in M$ one has, clearly, $\dim\mathfrak{hol}_{p}(\nabla^{C})\geq\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$. Because $\nabla^{C}$ is metric, its holonomy algebra sits in $\mathfrak{so}(T_{p}M)$. If the action were reducible with respect to $V_{1}\oplus V_{2}$, all holonomy endomorphisms would be block-diagonal, forcing $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\{0\}$. Hence nontrivial off-diagonal holonomy implies irreducibility.
4. Mixed Rank of a Degree-$3$ Class
Write the Künneth decomposition of $H^{3}(M)$ as:
$\big(H^{3}(M_{1})\otimes H^{0}(M_{2})\big)\ \oplus\ \big(H^{2}(M_{1})\otimes H^{1}(M_{2})\big)\ \oplus$ (4.1)
$\big(H^{1}(M_{1})\otimes H^{2}(M_{2})\big)\ \oplus\ \big(H^{0}(M_{1})\otimes H^{3}(M_{2})\big).$
Mixed components occur precisely in bidegrees $(2,1)$ and $(1,2)$.
Definition 4.1 (Mixed tensor rank).
Let $\omega]\in H^{3}(M;\mathbb{R})$. Denote by $\pi_{2,1}([\omega])\in H^{2}(M_{1})\otimes H^{1}(M_{2})$ and $\pi_{1,2}([\omega])\in H^{1}(M_{1})\otimes H^{2}(M_{2})$ its mixed Künneth projections. Define
$r_{2,1}=\min\Big\{r\,:\,\pi_{2,1}([\omega])=\sum_{i=1}^{r}\alpha_{i}\otimes\beta_{i},\ \alpha_{i}\in H^{2}(M_{1}),\ \beta_{i}\in H^{1}(M_{2})\Big\},$ (4.2)
and analogously $r_{1,2}$ for $\pi_{1,2}([\omega])$. The mixed rank is
$\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big):=r_{2,1}+r_{1,2}.$ (4.3)
For degree $3$, the tensor rank in each bidegree is well-defined and finite, and can be computed from any decomposition of the harmonic representative $\omega_{h}$ into wedge products of harmonic forms on the factors for the product metric $g$. Although $\omega_{h}$ depends on $g$, the ranks $r_{2,1},r_{1,2}$ are purely topological (indeed, the tensor rank of an element in $H^{2}(M_{1})\otimes H^{1}(M_{2})$ is defined as the minimal number of simple tensors in any decomposition, hence is invariant under change of metric within the product class).
5. Main Result and Proof
Lemma 5.1 (Existence of points of maximal rank for covariant derivatives).
Let $(M_{1},g_{1})$ be a compact oriented Riemannian manifold and let $\mathcal{H}^{2}(M_{1},g_{1})$ denote the space of harmonic $2$-forms. Given a subspace $U\subset\mathcal{H}^{2}(M_{1},g_{1})$ of dimension $r$ and its parallel part $U_{\parallel}:=U\cap\mathcal{P}_{2}(M_{1},g_{1})$ of dimension $s$, choose a basis $\alpha_{1},\dots,\alpha_{r}$ of $U$ such that $\alpha_{1},\dots,\alpha_{s}$ span $U_{\parallel}$. Then there exists a non‑empty open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
$\bigl(\nabla^{g_{1}}\alpha_{s+1},\dots,\nabla^{g_{1}}\alpha_{r}\bigr)$
are pointwise linearly independent; i.e.
$\mathcal{U}=\Bigl\{p\in M_{1}\;\Big|\;\operatorname{rank}\bigl[(\nabla^{g_{1}}\alpha_{i})(p)\bigr]_{i=s+1,\dots,r}=r-s\Bigr\}\neq\emptyset.$
Proof.
We proceed in three steps.
1. Reduction to a real‑analytic metric. Real‑analytic Riemannian metrics are dense in the space of smooth metrics on a compact manifold; this follows from the theorem of Greene–Jacobowitz on analytic approximation of smooth maps (see [4]). Hence we can choose a sequence of real‑analytic metrics $\{g_{1}^{(k)}\}_{k\in\mathbb{N}}$ converging to $g_{1}$ in the $C^{\infty}$ topology. The dimension of the space of parallel $2$-forms is upper‑semicontinuous with respect to the metric (see [3]); therefore, after passing to a subsequence, we may assume
$\dim\mathcal{P}_{2}\bigl(M_{1},g_{1}^{(k)}\bigr)\leq\dim\mathcal{P}_{2}(M_{1},g_{1})=s\qquad\text{for all }k.$
For each $k$ let $\alpha_{i}^{(k)}\in\mathcal{H}^{2}\bigl(M_{1},g_{1}^{(k)}\bigr)$ be the unique harmonic representative of the cohomology class $\alpha_{i}]$. By the stability theorem for harmonic forms under metric deformation (Warner [9], see also Kodaira–Spencer [7]), $\alpha_{i}^{(k)}\to\alpha_{i}$ in $C^{\infty}$. Consequently the covariant derivatives converge in $C^{1}$:
$\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}\longrightarrow\nabla^{g_{1}}\alpha_{i}\qquad(i=s+1,\dots,r).$
If we prove that for each $k$ there exists a non‑empty open set $\mathcal{U}^{(k)}\subset M_{1}$ where the $\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}$ are linearly independent, then the $C^{1}$ convergence guarantees that, for large enough $k$, the same linear independence holds on a non‑empty open set for the original metric $g_{1}$. Indeed, linear independence is equivalent to the non‑vanishing of at least one $(r-s)\times(r-s)$ minor of the matrix of components; each minor depends continuously on the metric and the forms, so a minor that is non‑zero on an open set for $g_{1}^{(k)}$ remains non‑zero on an open set for $g_{1}$ when $k$ is sufficiently large.
Thus it suffices to prove the lemma under the assumption that $g_{1}$ is real‑analytic.
2. The real‑analytic case. Assume now that $g_{1}$ is real‑analytic. Then every harmonic form is real‑analytic (by elliptic regularity for the Hodge–Laplacian, see [8]), and consequently the covariant derivatives $\nabla^{g_{1}}\alpha_{i}$ are real‑analytic sections of the vector bundle $E:=T^{*}M_{1}\otimes\Lambda^{2}T^{*}M_{1}$.
Consider the real‑analytic section
$\Phi:M_{1}\longrightarrow\Lambda^{r-s}E,\qquad\Phi(p):=(\nabla^{g_{1}}\alpha_{s+1})(p)\wedge\cdots\wedge(\nabla^{g_{1}}\alpha_{r})(p).$
The set $\mathcal{U}:=\{p\in M_{1}\mid\Phi(p)\neq 0\}$ is exactly the set where the covariant derivatives are linearly independent; because $\Phi$ is continuous, $\mathcal{U}$ is open. We must show that $\mathcal{U}\neq\emptyset$.
Suppose, for contradiction, that $\mathcal{U}=\emptyset$; i.e. $\Phi(p)=0$ for every $p\in M_{1}$. Then at each point the sections $\nabla^{g_{1}}\alpha_{s+1},\dots,\nabla^{g_{1}}\alpha_{r}$ are linearly dependent. Since the condition of rank $\leq d$ is given by the vanishing of all $(d+1)\times(d+1)$ minors, which are analytic functions, the locus where the rank is constant is an analytic subset. Let d<rsd<r-s be the minimal rank and let $p_{0}$ be a point where the rank equals $d$. There exists a connected open neighbourhood $V$ of $p_{0}$ on which the rank is constant and equal to $d$.
On $V$ we can therefore find real‑analytic functions $c_{1},\dots,c_{r-s}:V\to\mathbb{R}$, not all zero, such that
$\sum_{i=1}^{r-s}c_{i}(p)\bigl(\nabla^{g_{1}}\alpha_{s+i}\bigr)(p)=0\qquad\forall p\in V.$ (5.1)
Define the analytic $2$-form $\eta:=\sum_{i=1}^{r-s}c_{i}\alpha_{s+i}$ on $V$. From (5.1) we have $\nabla^{g_{1}}\eta=\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}+\sum_{i=1}^{r-s}c_{i}\nabla^{g_{1}}\alpha_{s+i}=\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}$.
Let $V\subset M_{1}$ be the non–empty open subset fixed above. Since the mixed cohomology class on $M_{1}\times M_{2}$ is assumed to be non–trivial, its harmonic representative has a non–vanishing component in $H^{2}(M_{1})\otimes H^{1}(M_{2})$. In particular, the projection onto $H^{2}(M_{1})$ is non–zero. Hence the harmonic forms $\alpha_{s+1},\dots,\alpha_{r}$ cannot all vanish identically on $M_{1}$. Let $W\subset M_{1}$ denote the complement of their common zero set. By real–analyticity, $W$ is a non–empty open subset of $M_{1}$.
Restricting to the non‑empty open set $V^{\prime}:=V\cap W$, we have on $V^{\prime}$
$\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}=0.$
Because the $\alpha_{s+i}(p)$ are linearly independent at every $p\in V^{\prime}$, this forces $dc_{i}=0$ on $V^{\prime}$ for each $i$. Thus the functions $c_{i}$ are constant on each connected component of $V^{\prime}$. By analyticity, they extend to constants on the whole connected open set $V$ (if two analytic functions coincide on an open set, they coincide everywhere on the connected component). Consequently, the $c_{i}$ are constant on $V$. Thus there exist constants $c_{1},\dots,c_{r-s}\in\mathbb{R}$, not all zero, such that
$\sum_{i=1}^{r-s}c_{i}\nabla^{g_{1}}\alpha_{s+i}=0\qquad\text{on }V.$
Define the global harmonic $2$-form
$\zeta:=\sum_{i=1}^{r-s}\bar{c}_{i}\alpha_{s+i}\in\mathcal{H}^{2}(M_{1}).$
Then $\nabla^{g_{1}}\zeta=0$ on $V$; i.e., $\zeta$ is parallel on the open set $V$. Because both $\zeta$ and the condition $\nabla^{g_{1}}\zeta=0$ are real-analytic, the unique-continuation principle [1] implies that $\nabla^{g_{1}}\zeta=0$ on the whole connected component of $M_{1}$ containing $V$, hence on all of $M_{1}$. Therefore $\zeta$ is a globally parallel $2$-form. But this contradicts the hypothesis that $\alpha_{s+1},\dots,\alpha_{r}$ are linearly independent modulo $\mathcal{P}_{2}(M_{1},g_{1})$. Indeed, $\zeta$ is a non-trivial linear combination of these forms that belongs to $\mathcal{P}_{2}(M_{1},g_{1})$. Hence our assumption that $\mathcal{U}=\emptyset$ is impossible, and we conclude that $\mathcal{U}$ is a non-empty open subset of $M_{1}$.
3. Returning to the original smooth metric. As explained in Step 1, the existence of a non‑empty open set of linear independence for a sequence of real‑analytic approximations $g_{1}^{(k)}$ together with the $C^{1}$ convergence of the covariant derivatives guarantees the existence of a non‑empty open set with the same property for the original smooth metric $g_{1}$. This completes the proof. ∎
Theorem 5.2 (Topological Lower Bound for Off-Diagonal Curvature).
Let $M=M_{1}\times M_{2}$ be an admissible compact oriented Riemannian product manifold (Definition 2.2) with product metric $g=g_{1}\oplus g_{2}$. Let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]$ is a non‑trivial mixed class.
Define the mixed factor subspaces
$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),$
$\mathcal{V}_{1,2}:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
where the spans are taken over any minimal harmonic decomposition of the mixed Künneth components of $\omega]$; in the admissible setting these spans are independent of the particular choice (Lemma 2.3).
Define the obstruction kernels
$\ker\Psi_{p}:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\quad\ker\widetilde{\Psi}_{p}:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
and set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$.
Then, for every $p\in M$,
$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}.$
The bound depends only on $\omega]_{\text{mixed}}$ and on the parallel-form strata $\mathcal{P}_{k}(M_{i})$; hence it is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.
Proof.
Let $\omega_{h}$ be the $g$-harmonic representative of $\omega]$. Decompose $\omega_{h}$ according to Künneth and Hodge on the factors. For the $(2,1)$-part, write
$\omega_{h}^{2,1}=\sum_{i=1}^{r_{2,1}}\alpha_{i}\wedge\beta_{i},$ (5.2)
with $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$ harmonic and chosen so that the number of terms is minimal. Similarly,
$\omega_{h}^{1,2}=\sum_{j=1}^{r_{1,2}}\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j},$ (5.3)
with $\tilde{\alpha}_{j}\in\mathcal{H}^{1}(M_{1})$ and $\tilde{\beta}_{j}\in\mathcal{H}^{2}(M_{2})$. The total number of simple mixed tensors equals $r:=r_{2,1}+r_{1,2}=\operatorname{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)$.
Set $r^{\sharp}:=r-\dim\mathcal{K}$. We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
Consider the curvature formula (2.5). The torsion $T$ decomposes as $T=\omega_{h}+\eta$, where $\omega_{h}$ is the harmonic representative of the mixed class $\omega]$ and $\eta=d\alpha+\delta\beta$ collects the exact and coexact parts.
Our aim is to show that the off‑diagonal holonomy algebra contains at least $r^{\sharp}$ linearly independent directions. These directions will be produced by the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ arising from the harmonic part $\omega_{h}$ alone. We must verify that the remaining contributions—the part of the linear term coming from $\eta$ and the quadratic term $Q_{T}$—cannot interfere with this lower bound.
Suppose that on some non‑empty open set $V\subset M$, the derivative $\nabla^{LC}\eta$ cancelled enough off‑diagonal directions from $\nabla^{LC}\omega_{h}$ to reduce the rank of $(\nabla^{LC}T)_{\mathrm{off}}$ below $r^{\sharp}$. This would impose linear relations among the covariant derivatives of the harmonic components $\alpha_{i}\wedge\beta_{i}$ on $V$. Applying the analytic‑approximation technique employed in Lemma 5.1 (see also Remark 5.6), we may work with real‑analytic metrics and forms. The relation then becomes an equality between real‑analytic sections, so by the unique continuation principle for analytic sections (Aronszajn [1]) it extends to the whole connected component containing $V$, hence globally. Since the tensors $\nabla^{LC}(\alpha_{i}\wedge\beta_{i})$ are pointwise linearly independent on the open set $\mathcal{U}$ provided by Lemma 5.1, evaluating the global relation at any point of $\mathcal{U}$ forces all coefficients to vanish, contradicting the existence of a non‑trivial cancellation. Hence, on a (possibly smaller) non‑empty open set, the linear term $(\nabla^{LC}T)_{\mathrm{off}}$ still exhibits at least $r^{\sharp}$ independent off‑diagonal directions originating from $\omega_{h}$. Regarding the quadratic term $Q_{T}$, observe that the $r^{\sharp}$ independent operators produced by the linear term belong to the off‑diagonal holonomy algebra by the Ambrose–Singer theorem [2]. The quadratic term contributes additional operators to the same algebra. Since the dimension of a vector space generated by a set of vectors is at least the rank of any linearly independent subset, adding further generators cannot reduce the dimension below $r^{\sharp}$. Consequently, the lower bound obtained from the linear term persists for the full curvature $R^{C}$. Thus, to establish $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ it suffices to analyse the linear term generated by the harmonic component $\omega_{h}$.
To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$. By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists a non‑empty open subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent. For any such $p$ we pick vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X_{i},Y_{i})|_{p}=1$ and $\beta_{i}(Z_{i})|_{p}=1$.
Define the linear map
$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i}),$ (5.4)
where the vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfy $\alpha_{i}(X_{i},Y_{i})=1$ and $\beta_{i}(Z_{i})=1$.
Claim 5.2.1 (Independence of choices). The subspace $\ker\Psi_{p}\subset\mathcal{V}_{2,1}$ is independent of the choice of normalizing vectors $\{X_{i},Y_{i},Z_{i}\}$ in  5.4.
Proof of Claim.
Writing $\omega_{h}^{2,1}=\sum_{i=1}^{r_{2,1}}\alpha_{i}\wedge\beta_{i}$, the product rule gives
$(\nabla^{LC}\omega_{h})^{\mathrm{off}}=\sum_{i=1}^{r_{2,1}}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i}).$
For vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$, $Z_{i}\in T_{p_{2}}M_{2}$ with $\alpha_{i}(X_{i},Y_{i})=1$, $\beta_{i}(Z_{i})=1$, the contraction $\iota_{Z_{i}}\iota_{Y_{i}}\iota_{X_{i}}$ extracts precisely the $i$-th component:
$(\nabla^{LC}\omega_{h})^{\mathrm{off}}(X_{i},Y_{i};Z_{i})=(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i}).$
Summing over $i$ with coefficients $c_{i}$ yields that $\Psi_{p}(\xi)=0$ is equivalent to the tensor equation
$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=0,$ (5.5)
which is manifestly independent of the vectors $\{X_{i},Y_{i},Z_{i}\}$. Hence $\ker\Psi_{p}$ is well-defined. ∎
Now, although the right-hand side of   5.4 depends on the choice of vectors, by Claim 5.2.1 the condition $\Psi_{p}(\xi)=0$ is independent of that choice.
If $\xi\in\ker\Psi_{p}$, then by definition
$\frac{1}{2}\sum_{i}c_{i}(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i})=0.$
Separating the bidegrees in   5.5 yields the two independent conditions
$\sum_{i=1}^{r_{2,1}}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0,\quad\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})=0.$ (5.6)
Since the harmonic forms $\beta_{i}$ are linearly independent in $\mathcal{H}^{1}(M_{2})$, the set of points where they are pointwise linearly independent is a non‑empty open subset of $M_{2}$ (its complement is the zero locus of certain smooth determinants). Denote this open set by $\mathcal{U}_{2}\subset M_{2}$.
Now fix any point $y_{0}\in\mathcal{U}_{2}$. For this fixed $y_{0}$, the first equation in (5.6) becomes:
$\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(x)\Bigr)\wedge\beta_{i}(y_{0})=0\quad\text{for all }x\in M_{1}.$
Since the covectors $\beta_{1}(y_{0}),\dots,\beta_{r_{2,1}}(y_{0})$ are linearly independent, this forces
$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(x)=0\quad\text{for all }x\in M_{1}.$
Analogously, because the harmonic forms $\alpha_{i}$ are linearly independent in $\mathcal{H}^{2}(M_{1})$, there exists a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ where they are pointwise linearly independent.
Fix $x_{0}\in\mathcal{U}_{1}$. The second equation in (5.6) then gives for all $y\in M_{2}$:
$\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}(x_{0})\Bigr)\wedge\nabla^{g_{2}}\beta_{i}(y)=0,$
which implies $\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}(y)=0$ for all $y\in M_{2}$.
Therefore,
$\eta_{1}:=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\eta_{2}:=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$ (5.7)
Conversely, if (5.7) holds, then $\nabla^{LC}\bigl(\sum_{i}c_{i}\alpha_{i}\wedge\beta_{i}\bigr)=0$, which forces $\Psi_{p}\bigl(\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\bigr)=0$. Therefore $\xi=\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}$ belongs to $\ker\Psi_{p}$ if and only if (5.7) is satisfied.
Now we identify this kernel with the intersection $\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr)$. If $\xi\in\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}\otimes\mathcal{P}_{1}\bigr)$, then $\xi$ can be written as a finite sum $\sum_{k}A_{k}\otimes B_{k}$ with $A_{k}\in\mathcal{P}_{2}$, $B_{k}\in\mathcal{P}_{1}$. For each term, $\nabla^{LC}(A_{k}\wedge B_{k})=0$ because $A_{k}$ and $B_{k}$ are parallel, hence $\Psi_{p}(A_{k}\otimes B_{k})=0$. By linearity, $\Psi_{p}(\xi)=0$, so $\xi\in\ker\Psi_{p}$.
For the converse, assume that $\xi=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$. Then (5.7) holds, i.e.
$\eta_{1}=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\eta_{2}=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$
Let $U=\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ and $V=\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$. Define
$U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1}),\qquad V_{\parallel}=V\cap\mathcal{P}_{1}(M_{2}),$
and set
$s=\dim U_{\parallel},\qquad t=\dim V_{\parallel}.$
Choose any basis $\{\alpha_{1},\dots,\alpha_{s}\}$ of $U_{\parallel}$ and complete it to a basis of $U$ by adding vectors $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ whose classes form a basis of $U/U_{\parallel}$ (hence linearly independent modulo $\mathcal{P}_{2}(M_{1})$).
Now apply Lemma 5.1 to the subspace $U\subset\mathcal{H}^{2}(M_{1})$ with its parallel part $U_{\parallel}$ (which has dimension $s$). The lemma guarantees a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}$, the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{r_{2,1}}$ are linearly independent.
Since $\eta_{1}$ is parallel, $\nabla^{g_{1}}\eta_{1}\equiv 0$; evaluating at $p_{1}$ we obtain
$\sum_{i=s+1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(p_{1})=0.$
The linear independence at $p_{1}$ forces $c_{i}=0$ for all $i=s+1,\dots,r_{2,1}$.
An identical argument applied to $M_{2}$ yields a point $p_{2}\in\mathcal{U}_{2}$ such that
$\sum_{i=t+1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}(p_{2})=0,$
and the linear independence of $\{\nabla^{g_{2}}\beta_{i}(p_{2})\}_{i=t+1}^{r_{2,1}}$ forces $c_{i}=0$ for all $i=t+1,\dots,r_{2,1}$.
Hence, the only possibly non-zero coefficients are those with $i\leq\min(s,t)$. For such indices, by construction $\alpha_{i}\in\mathcal{P}_{2}(M_{1})$ and $\beta_{i}\in\mathcal{P}_{1}(M_{2})$. Therefore,
$\xi=\sum_{i=1}^{\min(s,t)}c_{i}\alpha_{i}\otimes\beta_{i}\in\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2}).$
Since $\xi\in\mathcal{V}_{2,1}$ by hypothesis, we have shown
$\ker\Psi_{p}\subseteq\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr).$
The opposite inclusion has already been proved above. Thus
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr).$
An identical argument applied to the $(1,2)$-component yields
$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
Now consider the induced map on the quotient
$\overline{\Psi}_{p}:\mathcal{V}_{2,1}/\ker\Psi_{p}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\overline{\Psi}_{p}([\xi]):=\Psi_{p}(\xi).$
The map $\overline{\Psi}_{p}$ is well‑defined and linear. It is injective because $\overline{\Psi}_{p}([\xi])=0$ implies $\Psi_{p}(\xi)=0$, hence $\xi\in\ker\Psi_{p}$ and $\xi]=0$. Consequently, $\overline{\Psi}_{p}$ sends a basis of the quotient to a set of linearly independent off‑diagonal curvature operators, and we obtain
$\dim(\operatorname{im}\Psi_{p})=\dim\bigl(\mathcal{V}_{2,1}/\ker\Psi_{p}\bigr)=r_{2,1}-\dim\ker\Psi_{p}.$ (5.8)
An entirely analogous argument for the (1,2)(1,2)‑component gives
$\dim(\operatorname{im}\widetilde{\Psi}_{p})=r_{1,2}-\dim\ker\widetilde{\Psi}_{p}.$ (5.9)
Recall that $s=\dim(U_{\parallel})$ and $t=\dim(V_{\parallel})$. By Lemma 5.1 applied to the subspace $\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ with parallel part of dimension $s$, there exists a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}$ the covariant derivatives $\nabla^{g_{1}}\alpha_{i}(p_{1})$ (for $i=s+1,\dots,r_{2,1}$) are linearly independent. Similarly, applying Lemma 5.1 to $\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$ with parallel part of dimension $t$ yields a non‑empty open set $\mathcal{U}_{2}\subset M_{2}$ such that for every $p_{2}\in\mathcal{U}_{2}$ the covariant derivatives $\nabla^{g_{2}}\beta_{i}(p_{2})$ (for $i=t+1,\dots,r_{2,1}$) are linearly independent.
Consequently, the product set $\mathcal{U}:=\mathcal{U}_{1}\times\mathcal{U}_{2}$ is a non‑empty open subset of $M=M_{1}\times M_{2}$. For any point $p=(p_{1},p_{2})\in\mathcal{U}$ the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, i.e.
$\dim\operatorname{im}\Psi_{p}=r_{2,1}-\dim\ker\Psi_{p},\qquad\dim\operatorname{im}\widetilde{\Psi}_{p}=r_{1,2}-\dim\ker\widetilde{\Psi}_{p}.$
Hence, at every $p\in\mathcal{U}$, the off‑diagonal part of the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ generates at least
$\bigl(r_{2,1}-\dim\ker\Psi_{p}\bigr)+\bigl(r_{1,2}-\dim\ker\widetilde{\Psi}_{p}\bigr)=r-\dim\mathcal{K}=r^{\sharp}$
linearly independent curvature operators. By the Ambrose–Singer theorem [2], these operators belong to the holonomy algebra at $p$. By Definition 3.2 of the off‑diagonal holonomy algebra $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ as the Lie subalgebra generated by all off‑diagonal curvature components and their covariant derivatives, we therefore have
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}\qquad\text{for all }p\in\mathcal{U}.$
Now let $p_{0}$ be any point in $\mathcal{U}$. Because the manifold $M$ is connected, the holonomy algebras at different points are isomorphic via parallel transport along curves. In particular, the dimension of the off‑diagonal holonomy algebra is constant on $M$. Hence the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$, which holds at $p_{0}$, must hold at every point $p\in M$ and no density requirement is needed. This completes the proof of the lower bound.
The lower bound depends only on $\omega]$ and on the spaces of $g_{1}$‑parallel harmonic 22‑forms and 11‑forms, and of $g_{2}$‑parallel harmonic 11‑forms and 22‑forms; it is therefore invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these parallel‑form strata, hence under any product‑metric deformation that leaves $\mathcal{P}_{k}(M_{i})$ unchanged for $k=1,2$. ∎
Remark 5.3 (Necessity of the Non‑Parallel Torsion Condition).
Theorem 5.2 yields a non‑trivial lower bound only when the linear term
$\tfrac{1}{2}(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$
is non‑zero. If instead $\nabla^{\mathrm{LC}}T=0$, the curvature reduces to
$R^{C}=R^{\mathrm{LC}}+\tfrac{1}{4}Q_{T},$
and there is no guarantee that the purely quadratic off‑diagonal part $(Q_{T})_{\mathrm{off}}$ generates $r^{\sharp}$ independent directions.
Consider for instance the flat torus $M=T^{2}\times S^{1}$ with the constant volume form $T=dx\wedge dy\wedge dz$, which is harmonic and satisfies
$\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})=1.$
Since $T$ is parallel, $\nabla^{\mathrm{LC}}T=0$; hence every simple tensor in any minimal harmonic decomposition is completely parallel. In the notation of Theorem 5.2 we have
$\mathcal{V}_{2,1}=\operatorname{span}\{dx\wedge dy\otimes dz\},\qquad\mathcal{P}_{2}(T^{2})=\operatorname{span}\{dx\wedge dy\},\qquad\mathcal{P}_{1}(S^{1})=\operatorname{span}\{dz\},$
so that
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(T^{2})\otimes\mathcal{P}_{1}(S^{1})\bigr)=\mathcal{V}_{2,1},\qquad\ker\widetilde{\Psi}_{p}=0.$
Thus $\dim\mathcal{K}=1$ and consequently
$r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})-\dim\mathcal{K}=1-1=0.$
A direct computation gives $R^{C}=0$, hence
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})=0,\qquad r^{\sharp}=0.$
The bound $0\geq 0$ is saturated but trivial, illustrating that without the non‑parallel torsion condition the theorem cannot force irreducibility.
Remark 5.4 (Geometric Significance of Invariance).
The invariance of the lower bound $r^{\sharp}$ is a non‑trivial property that holds under the Assumption 2.2. In those cases, the bound depends only on:
  • the mixed class $\omega]_{\mathrm{mixed}}$,
  • the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ of the factor metrics.
Consequently, $r^{\sharp}$ is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata. This stability is relevant for manifolds with special holonomy where $\mathcal{P}_{k}(M_{i})$ are non‑trivial.
Remark 5.5 (Forced irreducibility and quantified off‑diagonal holonomy).
Since $\nabla^{C}$ is metric, any non‑trivial off‑diagonal holonomy implies that $\mathfrak{hol}_{p}(\nabla^{C})$ acts irreducibly on $T_{p}M$. Theorem 5.2 quantifies this “forced irreducibility” principle in terms of the mixed tensor rank, modulo the obstruction space
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},$
where, in the notation of the theorem,
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
Specifically, the dimension of the off‑diagonal holonomy subspace satisfies
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
and irreducibility is forced precisely when the right‑hand side is positive, i.e. when
$\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})>\dim\mathcal{K}.$
In other words, a non‑zero value of the corrected rank $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}$ guarantees that the holonomy representation cannot split according to the product decomposition $T_{p}M=V_{1}\oplus V_{2}$; the connection is geometrically “entangled” across the two factors.
Remark 5.6 (Analytic approximation as a tool of proof).
The proofs of Lemma 5.1 and Theorem 5.2 make essential use of real‑analytic approximations of the smooth metric. This technique does not restrict the generality of the results, because real‑analytic metrics are dense in the $C^{\infty}$ topology (Greene–Jacobowitz [4]) and the relevant quantities – the dimension of the space of parallel forms, the harmonic representatives of fixed cohomology classes, and the covariant derivatives of those harmonic forms – depend continuously on the metric in the appropriate topologies.
In Lemma 5.1 the reduction to an analytic metric allows us to apply the identity principle for analytic functions: if an analytic section of a vector bundle vanishes on an open set, it vanishes identically. This principle is used to show that a relation among the covariant derivatives of harmonic forms on an open set forces the existence of a parallel combination of the forms, contradicting the hypothesis of linear independence modulo parallel forms. For a smooth metric, we approximate by analytic metrics, prove the existence of a non‑empty open set of maximal rank for each approximation, and then pass to the limit using $C^{1}$ convergence of the covariant derivatives.
In Theorem 5.2 the analytic approximation is used only indirectly, through Lemma 5.1. Once the lemma guarantees a non‑empty open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows first on that open set, and then on the whole manifold because the dimension of the holonomy algebra is constant on connected components (Ambrose–Singer [2]).
Thus analyticity serves solely as a convenient tool to handle the delicate rank conditions in Lemma 5.1; the final statements are valid for arbitrary smooth product metrics and do not require any analyticity assumption on the metric or the torsion.
6. Illustrative Examples
The case $\Sigma_{g}\times T^{2}$ (Example of a non-trivial lower bound)
Fix $g\geq 2$ and endow $\Sigma_{g}$ with any hyperbolic metric and $T^{2}$ with the flat metric. Harmonic spaces:
$H^{1}(T^{2})=\operatorname{span}\{dx,dy\},\qquad H^{2}(\Sigma_{g})=\operatorname{span}\{\operatorname{vol}_{\Sigma}\},$
and choose a non-zero harmonic $1$-form $\alpha\in H^{1}(\Sigma_{g})$ (such an $\alpha$ exists because $b_{1}(\Sigma_{g})=2g>0$).
For the hyperbolic metric on $\Sigma_{g}$ we have
$\mathcal{P}_{1}(\Sigma_{g})=0,\qquad\mathcal{P}_{2}(\Sigma_{g})=\operatorname{span}\{\operatorname{vol}_{\Sigma}\}\cong\mathbb{R},$
since the volume form is harmonic and parallel.
For the flat metric on $T^{2}$ all harmonic forms are parallel, hence
$\mathcal{P}_{1}(T^{2})=H^{1}(T^{2})\cong\mathbb{R}^{2},\qquad\mathcal{P}_{2}(T^{2})=H^{2}(T^{2})\cong\mathbb{R}.$
Consider the mixed cohomology class
$\omega]=\underbrace{\operatorname{vol}_{\Sigma}\otimes dx}_{\in H^{2}(\Sigma_{g})\otimes H^{1}(T^{2})}\;+\;\underbrace{\alpha\otimes\operatorname{vol}_{T^{2}}}_{\in H^{1}(\Sigma_{g})\otimes H^{2}(T^{2})}\;\in\;H^{3}(\Sigma_{g}\times T^{2}).$
Both Künneth components are already simple tensors, therefore their minimal tensor ranks are
$r_{2,1}=1,\qquad r_{1,2}=1,$
and the mixed rank equals
$\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)=r_{2,1}+r_{1,2}=2.$
The obstruction kernels are computed as follows. For the $(2,1)$-part $\pi_{2,1}([\omega])=\operatorname{vol}_{\Sigma}\otimes dx$,
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr)=\operatorname{span}\{\operatorname{vol}_{\Sigma}\otimes dx\},$
because $\operatorname{vol}_{\Sigma}\in\mathcal{P}_{2}(\Sigma_{g})$ and $dx\in\mathcal{P}_{1}(T^{2})$. Hence $\dim\ker\Psi_{p}=1$.
For the $(1,2)$-part $\pi_{1,2}([\omega])=\alpha\otimes\operatorname{vol}_{T^{2}}$,
$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr)=\{0\},$
because $\mathcal{P}_{1}(M_{1})=\{0\}$ and $\mathcal{P}_{2}(M_{2})=\operatorname{span}\{\operatorname{vol}_{T^{2}}\}$, hence $\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})=\{0\}$. Hence $\dim\ker\widetilde{\Psi}_{p}=0$. Thus
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\dim\mathcal{K}=1+0=1,$
and the corrected rank is
$r^{\sharp}=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}=2-1=1.$
Theorem 5.2 therefore yields the non‑trivial lower bound
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;1.$
In particular, any cohomologically calibrated metric connection with non‑parallel torsion representing $\omega]$ necessarily has irreducible holonomy across the product splitting $\Sigma_{g}\times T^{2}$.
The case $T^{2}\times T^{2}$ (Example of a trivial lower bound)
Let $M=T^{2}\times T^{2}$ endowed with the flat product metric $g=g_{1}\oplus g_{2}$. Fix harmonic bases
$H^{1}(T^{2})=\operatorname{span}\{dx,dy\},\quad H^{2}(T^{2})=\operatorname{span}\{vol\}.$
For the flat metric we have
$\mathcal{P}_{1}(T^{2})=H^{1}(T^{2})\cong\mathbb{R}^{2},\qquad\mathcal{P}_{2}(T^{2})=H^{2}(T^{2})\cong\mathbb{R}.$
The mixed Künneth components are
$H^{2}(T^{2})\otimes H^{1}(T^{2})\cong\mathbb{R}^{2},\quad H^{1}(T^{2})\otimes H^{2}(T^{2})\cong\mathbb{R}^{2}.$
Consider the single cohomology class
$\omega]=\underbrace{vol\otimes dx}_{(2,1)}+\underbrace{dy\otimes vol}_{(1,2)}.$
Its minimal tensor-rank decomposition is
$\pi_{2,1}([\omega])=vol\otimes dx\quad(\text{rank }1),\qquad\pi_{1,2}([\omega])=dy\otimes vol\quad(\text{rank }1).$
Hence
$\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1+1=2.$
Since $vol\in\mathcal{P}_{2}$ and $dx,dy\in\mathcal{P}_{1}$, both simple tensors lie in the obstruction spaces
$\ker\Psi_{p}=\operatorname{span}\{vol\otimes dx\},\qquad\ker\tilde{\Psi}_{p}=\operatorname{span}\{dy\otimes vol\},$
so $\dim\mathcal{K}=2$. Theorem 5.2 gives
$\dim\mathfrak{hol}^{\text{off}}_{p}(\nabla^{C})\geq 2-2=0.$
7. Conclusion
We have shown that for admissible product manifolds $M=M_{1}\times M_{2}$—specifically, products of compact oriented surfaces or four‑dimensional $(3,1)$-splittings with rank‑one mixed classes—and for any cohomologically calibrated metric connection $\nabla^{C}$ with non‑parallel torsion, the off‑diagonal holonomy dimension admits a quantified lower bound
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\;\geq\;r^{\sharp}=r-\dim\mathcal{K},$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\operatorname{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the sum of the kernels of the linear obstruction maps. The bound is metric‑invariant under deformations preserving the parallel‑form strata $\mathcal{P}_{k}(M_{i})$. This provides the first computable obstruction to reducible holonomy on product manifolds within the admissible class.
The examples in Section 6 illustrate both non‑trivial and trivial cases, while Remark 2.4 shows by explicit construction that the bound fails to be well‑defined without the admissibility hypothesis, establishing sharpness of our assumptions.
When the torsion is purely harmonic and the parallel‑form strata are minimal, the bound is often saturated (e.g., on flat tori with parallel torsion). It is natural to conjecture that equality holds generically in this setting, while strict inequality may occur when the coexact/exact parts of $T$ generate additional off‑diagonal curvature or when the parallel‑form spaces are larger. A systematic classification of sharp cases remains an interesting open problem.
A further direction is to explore extensions beyond the admissible setting. Imposing Kruskal‑rank constraints on harmonic representatives could yield intrinsic mixed factor spaces even when $b_{2}(M_{i})>1$, potentially treating Calabi–Yau threefolds or $G_{2}\times S^{1}$ backgrounds. Such extensions would require new algebraic tools and are left for future investigation.
Another further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].
Compactness and Hodge theory. Compactness ensures the existence and uniqueness of harmonic representatives [9], and the vanishing of boundary terms in integrations by parts. Extending the results to noncompact manifolds would require an $L^{2}$ Hodge framework or alternative analytic hypotheses.